### v1

# Fanxu Meng1,2,3, Zhaohui Wang1, Muhan Zhang1,3

Abstract
As the parameters of large language models (LLMs) expand, the computational cost of fine-tuning the entire model becomes prohibitive. To address this challenge, we introduce a parameter-efficient fine-tuning (PEFT) method, Principal Singular values and Singular vectors Adaptation (PiSSA), which optimizes a significantly reduced parameter space while achieving or surpassing the performance of full-parameter fine-tuning. PiSSA is inspired by Intrinsic SAID, which suggests that pre-trained, over-parametrized models inhabit a space of low intrinsic dimension. Consequently, PiSSA represents a matrix $W\in\mathbb{R}^{m\times n}$ within the model by the product of two trainable matrices $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, where $r\ll\min(m,n)$, plus a residual matrix $W^{res}\in\mathbb{R}^{m\times n}$ for error correction. Singular value decomposition (SVD) is employed to factorize $W$, and the principal singular values and vectors of $W$ are utilized to initialize $A$ and $B$. The residual singular values and vectors initialize the residual matrix $W^{res}$, which keeps frozen during fine-tuning. Notably, PiSSA shares the same architecture with Low-Rank Adaptation (LoRA), which hypothesizes that changes in model parameters $\Delta W$ form a low-rank matrix. However, LoRA approximates $\Delta W$ through the product of two matrices, $A$, initialized with Gaussian noise, and $B$, initialized with zeros, while PiSSA initializes $A$ and $B$ with principal singular values and singular vectors of the original matrix $W$. Given that the principal singular values and vectors capture the essence of a low-rank matrix, PiSSA can better approximate the outcomes of full-parameter fine-tuning at the beginning by changing the essential parts while freezing the “noisy” parts. In comparison, LoRA freezes the original matrix and updates the “noise”. This distinction enables PiSSA to convergence much faster than LoRA and also achieve better performance in the end. On five common benchmarks, PiSSA outperforms LoRA on all of them using exactly the same setups except for a different initialization. On GSM8K, Mistral-7B fine-tuned with PiSSA achieves an accuracy of 72.86%, outperforming LoRA’s 67.7% by 5.16%. Due to the same architecture, PiSSA inherits many of LoRA’s advantages, such as parameter efficiency and compatibility with quantization. Leveraging a fast SVD method, the initialization of PiSSA takes only a few seconds, inducing negligible cost of switching LoRA to PiSSA.
1 Introduction
Fine-tuning large language models (LLMs) is a highly effective technique for boosting their capabilities in various tasks [1, 2, 3, 4], ensuring models to follow instructions [5, 6, 7], and instilling models with desirable behaviors while eliminating undesirable ones [8, 9]. However, the fine-tuning process for very large models is accompanied by prohibitive costs. For example, regular 16-bit fine-tuning of a LLaMA 65B parameter model requires over 780 GB of GPU memory [10], and the VRAM consumption for training GPT-3 175B reaches 1.2TB [11]. Consequently, various parameter-efficient fine-tuning (PEFT) [12, 13] methods have been proposed to reduce the number of parameters and memory usage required for fine-tuning. Due to the ability to maintain the performance of full fine-tuning without adding additional inference latency, Low-Rank Adaptation (LoRA) [11] has emerged as a popular PEFT method.
LoRA [11] hypothesizes that the modifications to parameter matrices during fine-tuning exhibit low-rank properties. As depicted in Figure 1(b), for a pre-trained weight matrix $W\in\mathbb{R}^{m\times n}$, LoRA substitutes the updates with a low-rank decomposition $\Delta W=AB$, where $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, and the rank $r\ll\text{min}(m,n)$. For $h=Wx$, the modified forward pass computation is expressed as follows:
$h=(W+\Delta W)x=(W+AB)x.$ (1)
A random Gaussian initialization is used for $A$ and zero for $B$, making $AB=0$ at the beginning of training, thereby the injection of adapters does not affect the model’s output initially. LoRA avoids the need to compute gradients or maintain the optimizer states for the original matrix, instead optimizing the injected, significantly smaller low-rank matrices. Thus, it could reduce the number of trainable parameters by 10,000 times and the GPU memory requirement by 3 times. Moreover, LoRA often achieves comparable or superior performance to full parameter fine-tuning, indicating that fine-tuning “parts” of the full parameters can be enough for downstream tasks. By integrating the quantization of pre-trained models, LoRA also enables reducing the average memory requirements by 16 times [10]. Meanwhile, the adapters are still allowed to utilize higher precision weights; thus, the quantization usually does not degrade the performance of full precision base models.
(a) Full Fine-tuning
(b) LoRA
(c) PiSSA
Figure 1: The comparison among Full Fine-tuning, training with LoRA, and PiSSA. In this visualization, blue modules represent parts of the model where parameters are frozen during training, while orange modules indicate components that require updates.
Inspired by the core idea of Intrinsic SAID [14, 15], our Principal Singular values and Singular vectors Adapter (PiSSA) diverges from LoRA and its successors by not focusing on approximating $\Delta W$, but $W$. In alignment with Intrinsic SAID, we assume that the matrix $W$ in pre-trained models exhibits low (intrinsic) rank characteristics. Thus, we apply singular value decomposition (SVD) to matrix $W$. Based on the magnitude of the singular values, we divide $W$ into two components: the principal matrix $W^{pri}$ and the residual matrix $W^{res}$. The principal matrix $W^{pri}$ is of low rank and can be represented by the product of $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, where $r\ll\min(m,n)$. $A$ and $B$ are initialized based on the principal singular values and vectors and are trainable, whereas $W^{res}$ is initialized with the product of the residual singular values and vectors and are frozen. As indicated by Equation (2), the integration of $AB$ (the PiSSA adaptor) with the residual model also preserves the capability of the pre-trained model in the beginning of fine-tuning:
$h=Wx=(W^{res}+W^{pri})x=(W^{res}+AB)x.$ (2)
Different from LoRA, a significantly less important part $W^{res}$ is kept frozen, while the essential part $W^{pri}$ is fully tunable from the beginning, which enables PiSSA to fit the training data faster and better.
2 Related Works
The vast complexity and computational needs of large language models (LLMs) with billions of parameters present significant hurdles in adapting them for specific downstream tasks. Parameter Efficient Fine-Tuning (PEFT) [12, 13] emerges as a compelling solution by minimizing the fine-tuning parameters and memory requirements while achieving comparable performance to full fine-tuning. PEFT encompasses strategies like partial fine-tuning, soft prompt fine-tuning, non-linear adapter fine-tuning, and linear adapter fine-tuning.
Partial fine-tuning selects certain model parts, e.g., bias and layer normalization [16, 17], or inserted trainable mask to the pre-trained weights [18, 19] and delta weights [20, 21, 22, 23], for fine-tuning. However, the performance of partial fine-tuning may be limited due to only specific layers’ parameters being fine-tuned. Soft prompt fine-tuning  [24, 25, 26, 27, 28, 29, 30] introduces optimizable vectors into the model’s input or hidden layers, adapting to specific tasks with minimal structural changes. Non-linear adapter fine-tuning  [31, 32, 33, 34, 35, 36, 37, 38, 39, 40] incorporates (in a non-linear manner) trainable (non-linear) modules to the models. Both soft prompt fine-tuning and non-linear adapter fine-tuning add complexity and inference costs despite their adaptability. Linear adapter fine-tuning [14, 15, 11, 41], particularly LoRA [11] and its derivative methods, involves adding trainable parameters to the self-attention and feed-forward layers, which are then linearly combined with the original weights. After fine-tuning, these adaptations can be reparametrized into the standard model structure, thus gaining widespread adoption due to their ability to maintain the model’s original architecture while enabling efficient fine-tuning.
LoRA [11] posits that weight adjustments during model adaptation exhibit a low "intrinsic rank". It learns the weight updates through a low-rank decomposition, keeping the original structure frozen and proving effective in parameter-efficient fine-tuning. Following LoRA, Low-rank Adjustment [42, 41, 43] learns the rank size needed for LoRA in each layer of the model. LoRA-guided Pretrained Weight Update [44, 45] updates the original weights of the model using the parameters from adapter layers, enhancing LoRA’s representational capacity. Quantization Adaptation [10, 46, 47] applies quantization to the base model, further reducing training consumption overhead.
Differing from LoRA’s approach of learning low-rank approximations of weight updates, our PiSSA method innovates by utilizing the low intrinsic dimension of pre-trained weight directly. It initializes updatable adapter layers with principal singular values and vectors from the model’s weights, preserving the essential capabilities while enabling effective fine-tuning. This approach, while different from LoRA’s in philosophy, shares most of the structural benefits of LoRA and can be combined with methods like Low-rank Adjustment for enhanced fine-tuning efficiency.
3 Method
[…]
$A$ $=U_{[:r]}\,\text{diag}(S_{[:r]}^{1/2})\in\mathbb{R}^{m\times r},$ (3)
$B$ $=\text{diag}(S_{[:r]}^{1/2})\,V_{[:r]}^{T}\in\mathbb{R}^{r\times n}.$ (4)
The residual singular values and vectors are used to build the residual matrix:
$W^{res}=U_{[r:]}\,\text{diag}(S_{[r:]})\,V_{[r:]}^{T}\in\mathbb{R}^{m\times n}.$ (5)
Given $W=W^{res}+AB$, the introduction of PiSSA in the beginning of training does not compromise the model’s capabilities. Considering $S_{[r:]}\approx 0$, thus rendering $W^{res}\approx 0$ and $W\approx AB$, the adapter $AB$ encapsulates the pre-trained model’s primary capabilities. Then, the fine-tuning of $AB$ mirrors the process of fine-tuning the entire model, despite using much less parameters. In contrast, LoRA initializes the adapter with Gaussian noise and zeros, thus the principal part of $W$ is not updatable in the early stage of the fine-tuning, possibly introducing much waste of gradient descent steps. Moreover, an inferior initialization might lead to suboptimal local minimum points found, causing worse generalization performance.
Since PiSSA shares the identical architecture with LoRA, it inherits most of LoRA’s benefits. These include but not limited to the capability of fine-tuning a model with a reduced number of trainable parameters, the quantization of the residual model to decrease memory consumption during forward propagation in training, and easy deployment. The adapter’s straightforward linear structure facilitates the integration of trainable matrices with the pre-trained weights upon deployment, thereby maintaining the original inference speed of a fully fine-tuned model. For storage efficiency, we can choose not to store the dense parameter matrix $\Delta W$, but to store the low-rank matrices, $\{A,B\}$ instead. When we want to utilize a fine-tuned PiSSA, similar to initializing one, we first perform singular value decomposition on the base model (which is very fast as will be shown in our experiments) and then assign $\{A,B\}$ to the adapter. Finally, one pre-trained model can accommodate multiple $\{A,B\}$, fine-tuned by diverse PiSSA procedures, which enables fast adapting of the pre-trained model to different downstream applications.
4 Experiments
We conduct comparative experiments between PiSSA, LoRA, and full parameter fine-tuning on LLaMA 2-7B [48], Mistral-7B-v0.1 [49], and Gemma-7B [50]. These comparing cover a wide range of tasks, including math [2, 51, 52], coding [4, 53, 54, 55], and instruction tuning [5, 7, 6]. The findings reveal that fine-tuning adapters initialized with principal singular values and vectors (PiSSA) outperform those initialized with intermediate or less significant ones. This evidences that direct fine-tuning of the model’s principal components leads to superior outcomes (Section 4.1). PiSSA demonstrates superior performance on downstream tasks (Section 4.3). Subsequently, we conduct experiments across various ranks, observing that PiSSA converges more swiftly and aligns more closely with the training data compared to LoRA, exhibiting robust superiority under the same trainable parameter configurations (Section 4.2). Finally, we employ subsequent experiments to demonstrate that the Fast SVD [56] technique helps our PiSSA achieve a balance between initialization speed and performance (Section 4.4).
In our experiments, we adopt the Alpaca [57] implementation strategy, using the AdamW optimizer with a batch size of 128, a learning rate of 2e-5, cosine annealing schedules, and a warmup ratio of 0.03, without any weight decay. We utilize the bf16 computation type, ensure lora_alpha is always equal to lora_r, set lora_dropout to 0, and incorporate the adapters into all linear layers of the base model. We train only on the first 100,000 conversions from the dataset for one epoch, and only using the response of the conversions to compute loss. The experiments in this section were conducted on the NVIDIA A800-SXM4(80G) GPU.
4.1 Verifying the importance of principal singular values and vectors
To investigate the influence of singular values and vectors of varying magnitudes on the fine-tuning performance, we initialize the adapters injected into LLaMA 2-7B, Mistral-7B-v0.1, and Gemma-7B with principal, medium, and minor singular values and vectors. These models are then fine-tuned on the MetaMathQA dataset [2] and evaluated against the GSM8K [51] and MATH datasets [52], with the outcomes depicted in Figures 2.
Figure 2: Initializing with principal, medium, and minor singular values and vectors, the training loss on the MetaMathQA and the accuracy on the GSM8K and MATH validation sets are reported, respectively, for three models.
The results highlight that initializing adapters with principal singular values and vectors consistently leads to reduced training loss and enhanced accuracy on both the GSM8K and MATH validation datasets across all three models. This underscores the efficacy of our strategy in fine-tuning the model parameters based on the principal singular values.
4.2 Evaluating PiSSA and LoRA Performance Across Diverse Ranks
This section explores the impact of incrementally increasing the rank of PiSSA and LoRA from 1 to 128, aiming to determine the optimal number of trainable parameters needed to achieve performance comparable to full parameter training. The training is conducted using the MetaMathQA dataset, while the validation is performed on the GSM8K and MATH datasets. The outcomes of these experiments are depicted in Figure 3, with additional results presented in Figure 4 in the Appendix.
(a) Variation of loss with respect to rank 1 throughout the training phase. Additional ranks are depicted in Figure 4 in the Appendix.
(b) Final training loss across different ranks.
(c) Rank-wise performance evaluated using pass@1 on the GSM8K dataset.
(d) Rank-wise performance evaluated using pass@1 on the MATH dataset.
Figure 3: Fine-tuning LLaMA 2-7B, Mistral-7B-v0.1, and Gemma-7B on the MetaMathQA dataset: A comparison of full parameter fine-tuning (indicated by a dashed line), LoRA (in blue), and PiSSA (in orange).
From Figure 3(a), it is observed that at the beginning of training, PiSSA quickly starts to converge, whereas LoRA experiences a period where the loss not only fails to decrease but even slightly increases. Furthermore, throughout the entire training process, the loss of PiSSA consistently remains at a lower level compared to LoRA. Figure 3(b) shows the final loss on the training set for models trained with ranks varying from 1 to 128. The data indicate that, compared to LoRA, PiSSA achieves a level of fitting to the training set similar to that of full-parameter fine-tuning with fewer trainable parameters. In Figures 3(c) and 3(d), we compare the performance of fine-tuned models on the GSM8K and MATH validation sets under various ranks, finding that PiSSA outperforms LoRA with the same amount of trainable parameters. Furthermore, as the rank increases, PiSSA quickly reaches and surpasses the performance of full-parameter fine-tuning. Notably, when the training loss of LLaMA 2-7B with a rank of 128 PiSSA equals that of full-parameter fine-tuning, it surpasses full-parameter fine-tuning in terms of generalization on downstream tasks. This validates our approach of using principal singular values and vectors to replace full-parameter model fine-tuning, ignoring components of lower rank, resulting in a model with improved robustness.
4.3 Comparative Analysis of PiSSA, LoRA, and Full Fine-Tuning Across Various Tasks
Table 1 presents a comparative evaluation of fine-tuning strategies using PiSSA, LoRA, and full parameter fine-tuning. In these experiments, we fine-tune LLaMA 2-7B, Mistral-7B-v0.1, and Gemma-7B on the MetaMathQA dataset [2] to assess their mathematical problem-solving capabilities on the GSM8K [51] and MATH [2] validation sets. Additionally, the models are fine-tuned on the CodeFeedback dataset [53] and their coding proficiency is evaluated on the HumanEval [54] and MBPP [55] datasets. Furthermore, the models are trained on the WizardLM-Evol-Instruct dataset [7] and tested for their conversational abilities on the MT-Bench dataset [6].
Table 1: Performance of different models on various tasks
Model Strategy Trainable GSM8K MATH HumanEval MBPP MT-Bench
Parameters
LLaMA 2-7B Full Fine-tuning 6738M 49.05 7.22 21.34 35.59 4.91
LoRA 320M 42.3 5.5 18.29 35.34 4.58
PiSSA 320M 53.07 7.44 21.95 37.09 4.87
Mistral-7B Full Fine-tuning 7242M 67.02 18.6 45.12 51.38 4.95
LoRA 168M 67.7 19.68 43.9 58.39 4.9
PiSSA 168M 72.86 21.54 46.95 62.66 5.34
Gemma-7B Full Fine-tuning 8538M 71.34 22.74 46.95 55.64 5.4
LoRA 200M 74.9 31.28 53.66 65.41 4.98
PiSSA 200M 77.94 31.94 54.27 66.17 5.64
As can be seen from Table 1, across all models and tasks, fine-tuning with PiSSA surpasses the effects of fine-tuning with LoRA, as well as most of full fine-tuning. In the future, we will extend the comparing to more models and tasks.
Table 2: Comparation between SVD and Fast SVD in terms of initialization time, error and training loss.
Metric Niter Rank
1 2 4 8 16 32 64 128
Initialization Time 1 5.05 8.75 5.07 8.42 5.55 8.47 6.80 11.89
2 4.38 4.71 4.79 4.84 5.06 5.79 7.70 16.75
4 5.16 4.73 5.09 5.16 5.60 7.01 7.90 11.41
8 4.72 5.11 5.14 5.40 5.94 7.80 10.09 14.81
16 6.24 6.57 6.80 7.04 7.66 9.99 14.59 22.67
$\infty$ 434.92 434.15 434.30 435.42 435.25 437.22 434.48 435.84
Initialization Error 1 1.30E-3 1.33E-3 1.55E-3 1.9E-3 1.98E-3 1.97E-3 2.00E-3 1.93E-3
2 5.84E-4 1.25E-3 1.45E-3 1.43E-3 1.48E-3 1.55E-3 1.48E-3 1.33E-3
4 6.01E-4 8.75E-4 6.75E-4 1.10E-3 1.05E-3 1.03E-3 1.08E-3 9.75E-4
8 1.26E-4 2.34E-4 5.25E-4 7.25E-4 5.75E-4 8.25E-4 8.25E-4 7.75E-4
16 7.93E-5 2.25E-4 1.28E-4 6.50E-4 4.25E-4 6.50E-4 6.00E-4 4.75E-4
$\infty$
Training Loss 1 0.3629 0.3420 0.3237 0.3044 0.2855 0.2657 0.2468 0.2301
2 0.3467 0.3337 0.3172 0.2984 0.2795 0.2610 0.2435 0.2282
4 0.3445 0.3294 0.3134 0.2958 0.2761 0.2581 0.2414 0.2271
8 0.3425 0.3279 0.3122 0.2950 0.2753 0.2571 0.2406 0.2267
16 0.3413 0.3275 0.3116 0.2946 0.2762 0.2565 0.2405 0.2266
$\infty$ 0.3412 0.3269 0.3116 0.2945 0.2762 0.2564 0.2403 0.2264
4.4 Fast Singular Value Decomposition
In order to speed up the decomposition of the pre-trained matrix $W$, we adopted the algorithm proposed by Halko et.al [56] (denoted as Fast SVD), which introduces randomness to achieve an approximate matrix decomposition. We compare the initialization time, error, and training loss between SVD and Fast SVD, with the results shown in Table 4.3. Initialization time refers to the computation time taken to decompose the pre-trained parameter matrix $W$, measured in seconds. Initialization error indicates the magnitude of the discrepancy introduced by Fast SVD compared to SVD after decomposing the matrix. Specifically, the error is the sum of the absolute differences between the matrices decomposed by original SVD and Fast SVD. For the error, we report the results of the self-attention module in the table. Loss refers to the loss value at the end of training. In Fast SVD, the parameter niter refers to the number of subspace iterations to conduct. A larger niter leads to increased decomposition time but results in smaller decomposition error. The symbol $\infty$ represents the experimental results with the SVD method.
It can be observed that the computation time of the SVD is tens of times that of Fast SVD. In addition, SVD exhibits consistently high time consumption with minimal variation as the rank increases, while Fast SVD, although experiencing a slight increase in computation time with higher ranks, remains significantly lower compared to SVD throughout. As the rank increases, the initialization error initially rises gradually, with a slight decrease observed when the rank reaches 128. And at the same rank, increasing the niter in Fast SVD leads to a gradual reduction in error. For training loss, we observed that as the rank increases, the training loss decreases gradually. At the same rank, with the increase of niter, the training loss of models initialized based on Fast SVD approaches that of models initialized based on SVD.
5 Conclusion
This paper introduces a parameter-efficient fine-tuning (PEFT) technique in which singular value decomposition is applied to the weight matrix of pre-trained models to extract principal components. These components are then used to initialize an adapter named PiSSA. Fine-tuning PiSSA can closely replicate the effects of fine-tuning the complete model in the beginning, while maintaining a good parameter efficiency. Through extensive experiments, we show that PiSSA’s fine-tuning performance significantly surpasses that of LoRA. PiSSA provides a novel direction for research in PEFT by identifying and fine-tuning the principal components within the model, analogous to slicing and re-baking the richest slice of a pizza. As PiSSA shares the same architecture as LoRA, it is easy to use and provides an efficient alternative initialization method. There remain several questions to be verified in the future: 1) Can PiSSA achieve significant improvements on a broader range of tasks and larger models? 2) When the iteration steps of LoRA are sufficiently long (adequately fitting the data), can it match the performance of PiSSA? 3) Can combining the successors of LoRA with PiSSA lead to further enhancement? 4) How to give a theoretical explanation for the advantages of PiSSA over LoRA? We are actively exploring these questions. Nevertheless, we are excited to see the huge potential of PiSSA already demonstrated in existing experiments, and look forward to more tests and suggestions from the community.

### v2

@@
 Due to the same architecture, PiSSA inherits many of LoRA’s advantages, such as parameter efficiency and compatibility with quantization.
-Leveraging a fast SVD method, the initialization of PiSSA takes only a few seconds, inducing negligible cost of switching LoRA to PiSSA.
+Furthermore, PiSSA reduces the 4-bit quantization error in LLaMA 2-7B by 18.97%, resulting in a substantial improvement in fine-tuning performance.
+On the GSM8K benchmark, PiSSA achieves an accuracy of 49.13%, surpassing the performances of QLoRA at 39.8% and LoftQ at 40.71%.
+Leveraging a fast SVD technique, the initialization of PiSSA takes only a few seconds, inducing negligible cost of switching LoRA to PiSSA.
 1 Introduction
@@
 By integrating the quantization of pre-trained models, LoRA also enables reducing the average memory requirements by 16 times [10].
-Meanwhile, the adapters are still allowed to utilize higher precision weights; thus, the quantization usually does not degrade the performance of full precision base models.
+Meanwhile, the adapters are still allowed to utilize higher precision weights; thus, the quantization usually does not degrade the performance of full-precision base models.
 (a) Full Fine-tuning
@@
 Different from LoRA, a significantly less important part $W^{res}$ is kept frozen, while the essential part $W^{pri}$ is fully tunable from the beginning, which enables PiSSA to fit the training data faster and better.
+Since the high-frequency principal component $W^{pri}$ is preserved in the adapter, when fine-tuning PiSSA incorporated with quantization, PiSSA significantly reduced quantization errors in the initial stages compared to using QLoRA.
 2 Related Works
@@
 It initializes updatable adapter layers with principal singular values and vectors from the model’s weights, preserving the essential capabilities while enabling effective fine-tuning.
-This approach, while different from LoRA’s in philosophy, shares most of the structural benefits of LoRA and can be combined with methods like Low-rank Adjustment for enhanced fine-tuning efficiency.
+This approach, while different from LoRA’s in philosophy, shares most of the structural benefits of LoRA and can be combined with methods like quantization the residual model for reducing computation cost.
 3 Method
@@
 Given $W=W^{res}+AB$, the introduction of PiSSA in the beginning of training does not compromise the model’s capabilities.
-Considering $S_{[r:]}\approx 0$, thus rendering $W^{res}\approx 0$ and $W\approx AB$, the adapter $AB$ encapsulates the pre-trained model’s primary capabilities.
+Considering $S_{[:r]}\gg S_{[r:]}$, thus rendering $W^{pri}\gg W^{res}$ and $W\approx AB$, the adapter $AB$ encapsulates the pre-trained model’s primary capabilities.
 Then, the fine-tuning of $AB$ mirrors the process of fine-tuning the entire model, despite using much less parameters.
@@
 Since PiSSA shares the identical architecture with LoRA, it inherits most of LoRA’s benefits.
-These include but not limited to the capability of fine-tuning a model with a reduced number of trainable parameters, the quantization of the residual model to decrease memory consumption during forward propagation in training, and easy deployment.
+These include, but are not limited to, the capability of fine-tuning a model with a reduced number of trainable parameters, quantizing the residual model to decrease memory consumption during forward propagation in training, and easy deployment, during which process PiSSA can significantly reduce the initial quantization error in the initial stages.
 The adapter’s straightforward linear structure facilitates the integration of trainable matrices with the pre-trained weights upon deployment, thereby maintaining the original inference speed of a fully fine-tuned model.
-For storage efficiency, we can choose not to store the dense parameter matrix $\Delta W$, but to store the low-rank matrices, $\{A,B\}$ instead.
-When we want to utilize a fine-tuned PiSSA, similar to initializing one, we first perform singular value decomposition on the base model (which is very fast as will be shown in our experiments) and then assign $\{A,B\}$ to the adapter.
-Finally, one pre-trained model can accommodate multiple $\{A,B\}$, fine-tuned by diverse PiSSA procedures, which enables fast adapting of the pre-trained model to different downstream applications.
+For storage efficiency, we can choose not to store the dense parameter matrix $\Delta W$, but to store the low-rank matrices, $\Delta AB$ instead.
+As shown in Appendix A, leveraging solely the $\Delta AB$ facilitates their seamless integration with the original pre-trained models.
+Finally, one pre-trained model can accommodate multiple $\Delta AB$, fine-tuned by diverse PiSSA procedures, which enables fast adapting of the pre-trained model to different downstream applications.
 4 Experiments
-We conduct comparative experiments between PiSSA, LoRA, and full parameter fine-tuning on LLaMA 2-7B [48], Mistral-7B-v0.1 [49], and Gemma-7B [50].
-These comparing cover a wide range of tasks, including math [2, 51, 52], coding [4, 53, 54, 55], and instruction tuning [5, 7, 6].
-The findings reveal that fine-tuning adapters initialized with principal singular values and vectors (PiSSA) outperform those initialized with intermediate or less significant ones.
-This evidences that direct fine-tuning of the model’s principal components leads to superior outcomes (Section 4.1).
-PiSSA demonstrates superior performance on downstream tasks (Section 4.3).
-Subsequently, we conduct experiments across various ranks, observing that PiSSA converges more swiftly and aligns more closely with the training data compared to LoRA, exhibiting robust superiority under the same trainable parameter configurations (Section 4.2).
-Finally, we employ subsequent experiments to demonstrate that the Fast SVD [56] technique helps our PiSSA achieve a balance between initialization speed and performance (Section 4.4).
+We conducted comparative experiments among PiSSA, LoRA, and full parameter fine-tuning on three models: LLaMA 2-7B [48], Mistral-7B-v0.1 [49], and Gemma-7B [50].
+Our comparison spanned a diverse range of tasks, encompassing mathematics [2, 51, 52], coding [4, 53, 54, 55], and instruction tuning [5, 7, 6].
+PiSSA exhibited superior performance across these downstream tasks (Section 4.1).
+When combined with quantization techniques, PiSSA effectively reduces quantization errors.
+Compared to QLoRA [10] and LoftQ [47], PiSSA exhibits superior performance both during initialization and fine-tuning stages (Section 4.2).
+The ablation studies, which varied across different ranks, showed that PiSSA converged more swiftly and aligned more closely with the training data than LoRA, demonstrating robust superiority under equivalent trainable parameter configurations (Section 4.3).
+Further experiments demonstrated that employing the Fast SVD technique [56] allowed PiSSA to achieve a balance between initialization speed and performance (Section 4.4).
+Lastly, adapters initialized with principal singular values and vectors (PiSSA) were compared to those initialized with intermediate or less significant ones, revealing that direct fine-tuning of the model’s principal components yielded better outcomes (Section 4.5).
 In our experiments, we adopt the Alpaca [57] implementation strategy, using the AdamW optimizer with a batch size of 128, a learning rate of 2e-5, cosine annealing schedules, and a warmup ratio of 0.03, without any weight decay.
@@
 The experiments in this section were conducted on the NVIDIA A800-SXM4(80G) GPU.
-4.1 Verifying the importance of principal singular values and vectors
-To investigate the influence of singular values and vectors of varying magnitudes on the fine-tuning performance, we initialize the adapters injected into LLaMA 2-7B, Mistral-7B-v0.1, and Gemma-7B with principal, medium, and minor singular values and vectors.
-These models are then fine-tuned on the MetaMathQA dataset [2] and evaluated against the GSM8K [51] and MATH datasets [52], with the outcomes depicted in Figures 2.
-Figure 2: Initializing with principal, medium, and minor singular values and vectors, the training loss on the MetaMathQA and the accuracy on the GSM8K and MATH validation sets are reported, respectively, for three models.
-The results highlight that initializing adapters with principal singular values and vectors consistently leads to reduced training loss and enhanced accuracy on both the GSM8K and MATH validation datasets across all three models.
-This underscores the efficacy of our strategy in fine-tuning the model parameters based on the principal singular values.
-4.2 Evaluating PiSSA and LoRA Performance Across Diverse Ranks
-This section explores the impact of incrementally increasing the rank of PiSSA and LoRA from 1 to 128, aiming to determine the optimal number of trainable parameters needed to achieve performance comparable to full parameter training.
-The training is conducted using the MetaMathQA dataset, while the validation is performed on the GSM8K and MATH datasets.
-The outcomes of these experiments are depicted in Figure 3, with additional results presented in Figure 4 in the Appendix.
-(a) Variation of loss with respect to rank 1 throughout the training phase. Additional ranks are depicted in Figure 4 in the Appendix.
-(b) Final training loss across different ranks.
-(c) Rank-wise performance evaluated using pass@1 on the GSM8K dataset.
-(d) Rank-wise performance evaluated using pass@1 on the MATH dataset.
-Figure 3: Fine-tuning LLaMA 2-7B, Mistral-7B-v0.1, and Gemma-7B on the MetaMathQA dataset: A comparison of full parameter fine-tuning (indicated by a dashed line), LoRA (in blue), and PiSSA (in orange).
-From Figure 3(a), it is observed that at the beginning of training, PiSSA quickly starts to converge, whereas LoRA experiences a period where the loss not only fails to decrease but even slightly increases.
-Furthermore, throughout the entire training process, the loss of PiSSA consistently remains at a lower level compared to LoRA.
-Figure 3(b) shows the final loss on the training set for models trained with ranks varying from 1 to 128.
-The data indicate that, compared to LoRA, PiSSA achieves a level of fitting to the training set similar to that of full-parameter fine-tuning with fewer trainable parameters.
-In Figures 3(c) and 3(d), we compare the performance of fine-tuned models on the GSM8K and MATH validation sets under various ranks, finding that PiSSA outperforms LoRA with the same amount of trainable parameters.
-Furthermore, as the rank increases, PiSSA quickly reaches and surpasses the performance of full-parameter fine-tuning.
-Notably, when the training loss of LLaMA 2-7B with a rank of 128 PiSSA equals that of full-parameter fine-tuning, it surpasses full-parameter fine-tuning in terms of generalization on downstream tasks.
-This validates our approach of using principal singular values and vectors to replace full-parameter model fine-tuning, ignoring components of lower rank, resulting in a model with improved robustness.
 4.1 Comparative Analysis of PiSSA, LoRA, and Full Fine-Tuning Across Various Tasks
@@
 In the future, we will extend the comparing to more models and tasks.
+4.2 Comparative Analysis of QLoRA, LoftQ, and PiSSA in Quantization Fine-tuning
+In this section, we compare the performance of QLoRA, LoftQ, and PiSSA in fine-tuning with quantization.
+All base models are quantized to ‘nf4’ precision and frozen, while the adapters are trained using ‘bfloat16’ precision.
+We begin by comparing the quantization errors of the three methods.
+For each method, the quantized baseline model is combined with its adapter, and the difference between the original weights and this combination serves as the error matrix.
+We measure the magnitude of the errors by the average nuclear norm of all error matrices.
+Using the quantization error of the original model without any adapter as a baseline, the errors for QLoRA, LoftQ, and PiSSA are divided by this baseline to express them as error rates.
+As QLoRA’s adapter employs Gaussian-zero initialization, the error rate does not decrease; LoftQ can eliminate quantization errors through multiple singular value decompositions, but for consistency with PiSSA, only one decomposition is performed.
+PiSSA requires a full-precision singular value decomposition, storing the principal singular values and vectors in the adapter before quantizing the residual model.
+We train LLaMA 2-7B on MetaMathQA and validate the effectiveness of the three methods on GSM8K and MATH datasets.
+The reductions in error and the training results are displayed in Figure 2.
+Figure 2: Comparing the quantization error, the fine-tuning loss on the MetaMathQA and the accuracy on the GSM8K and MATH validation sets.
+The results, shown in Figure 2, reveal that PiSSA significantly outperforms LoftQ in reducing errors, achieving a 19% reduction at rank 128, as opposed to LoftQ’s 8%.
+LoftQ’s fine-tuning improvements over QLoRA are not significant, and in certain settings, it even underperforms compared to QLoRA.
+PiSSA effectively reduces quantization errors, and the training of principal singular values and vectors additionally provides benefits, thus substantially enhancing training outcomes in comparison to QLoRA and LoftQ.
+4.3 Evaluating PiSSA and LoRA Performance Across Diverse Ranks
+This section explores the impact of incrementally increasing the rank of PiSSA and LoRA from 1 to 128, aiming to determine the optimal number of trainable parameters needed to achieve performance comparable to full parameter training.
+The training is conducted using the MetaMathQA dataset, while the validation is performed on the GSM8K and MATH datasets.
+The outcomes of these experiments are depicted in Figure 3, with additional results presented in Figure 5 in the Appendix.
+(a) Variation of loss with respect to rank 1 throughout the training phase. Additional ranks are depicted in Figure 5 in the Appendix.
+(b) Final training loss across different ranks.
+(c) Rank-wise performance evaluated using pass@1 on the GSM8K dataset.
+(d) Rank-wise performance evaluated using pass@1 on the MATH dataset.
+Figure 3: Fine-tuning LLaMA 2-7B, Mistral-7B-v0.1, and Gemma-7B on the MetaMathQA dataset: A comparison of full parameter fine-tuning (indicated by a dashed line), LoRA (in blue), and PiSSA (in orange).
+From Figure 3(a), it is observed that at the beginning of training, PiSSA quickly starts to converge, whereas LoRA experiences a period where the loss not only fails to decrease but even slightly increases.
+Furthermore, throughout the entire training process, the loss of PiSSA consistently remains at a lower level compared to LoRA.
+Figure 3(b) shows the final loss on the training set for models trained with ranks varying from 1 to 128.
+The data indicate that, compared to LoRA, PiSSA achieves a level of fitting to the training set similar to that of full-parameter fine-tuning with fewer trainable parameters.
+In Figures 3(c) and 3(d), we compare the performance of fine-tuned models on the GSM8K and MATH validation sets under various ranks, finding that PiSSA outperforms LoRA with the same amount of trainable parameters.
+Furthermore, as the rank increases, PiSSA quickly reaches and surpasses the performance of full-parameter fine-tuning.
+Notably, when the training loss of LLaMA 2-7B with a rank of 128 PiSSA equals that of full-parameter fine-tuning, it surpasses full-parameter fine-tuning in terms of generalization on downstream tasks.
+This validates our approach of using principal singular values and vectors to replace full-parameter model fine-tuning, ignoring components of lower rank, resulting in a model with improved robustness.
+4.4 Fast Singular Value Decomposition
+In order to speed up the decomposition of the pre-trained matrix $W$, we adopted the algorithm proposed by Halko et.al [56] (denoted as Fast SVD), which introduces randomness to achieve an approximate matrix decomposition.
+We compare the initialization time, error, and training loss between SVD and Fast SVD, with the results shown in Table 4.4.
+Initialization time refers to the computation time taken to decompose the pre-trained parameter matrix $W$, measured in seconds.
+Initialization error indicates the magnitude of the discrepancy introduced by Fast SVD compared to SVD after decomposing the matrix.
+Specifically, the error is the sum of the absolute differences between the matrices decomposed by original SVD and Fast SVD.
+For the error, we report the results of the self-attention module in the table.
+Loss refers to the loss value at the end of training.
+In Fast SVD, the parameter niter refers to the number of subspace iterations to conduct.
+A larger niter leads to increased decomposition time but results in smaller decomposition error.
+The symbol $\infty$ represents the experimental results with the SVD method.
 Table 2: Comparation between SVD and Fast SVD in terms of initialization time, error and training loss.
@@
 $\infty$ 0.3412 0.3269 0.3116 0.2945 0.2762 0.2564 0.2403 0.2264
-4.4 Fast Singular Value Decomposition
-In order to speed up the decomposition of the pre-trained matrix $W$, we adopted the algorithm proposed by Halko et.al [56] (denoted as Fast SVD), which introduces randomness to achieve an approximate matrix decomposition.
-We compare the initialization time, error, and training loss between SVD and Fast SVD, with the results shown in Table 4.3.
-Initialization time refers to the computation time taken to decompose the pre-trained parameter matrix $W$, measured in seconds.
-Initialization error indicates the magnitude of the discrepancy introduced by Fast SVD compared to SVD after decomposing the matrix.
-Specifically, the error is the sum of the absolute differences between the matrices decomposed by original SVD and Fast SVD.
-For the error, we report the results of the self-attention module in the table.
-Loss refers to the loss value at the end of training.
-In Fast SVD, the parameter niter refers to the number of subspace iterations to conduct.
-A larger niter leads to increased decomposition time but results in smaller decomposition error.
-The symbol $\infty$ represents the experimental results with the SVD method.
 It can be observed that the computation time of the SVD is tens of times that of Fast SVD.
@@
 At the same rank, with the increase of niter, the training loss of models initialized based on Fast SVD approaches that of models initialized based on SVD.
+4.5 Verifying the importance of principal singular values and vectors
+To investigate the influence of singular values and vectors of varying magnitudes on the fine-tuning performance, we initialize the adapters injected into LLaMA 2-7B, Mistral-7B-v0.1, and Gemma-7B with principal, medium, and minor singular values and vectors.
+These models are then fine-tuned on the MetaMathQA dataset [2] and evaluated against the GSM8K [51] and MATH datasets [52], with the outcomes depicted in Figures 4.
+Figure 4: Initializing with principal, medium, and minor singular values and vectors, the training loss on the MetaMathQA and the accuracy on the GSM8K and MATH validation sets are reported, respectively, for three models.
+The results highlight that initializing adapters with principal singular values and vectors consistently leads to reduced training loss and enhanced accuracy on both the GSM8K and MATH validation datasets across all three models.
+This underscores the efficacy of our strategy in fine-tuning the model parameters based on the principal singular values.
 5 Conclusion
@@
 Nevertheless, we are excited to see the huge potential of PiSSA already demonstrated in existing experiments, and look forward to more tests and suggestions from the community.
+Appendix A Equivalently Convert PiSSA into LoRA
+The advantage of PiSSA lies in its ability to significantly enhance training outcomes during the fine-tuning phase.
+After training, it allows for the direct sharing of the trained matrices $A$ and $B$.
+However, users need to perform singular value decomposition on the original model, which requires additional time.
+When employing fast singular value decomposition, there can be slight inaccuracies.
+More importantly, utilizing PiSSA for inference necessitates altering the parameters of the original model, which can be inconvenient when using multiple adapters, especially when some adapters might be disabled or activated.
+Therefore, we recommend converting the trained PiSSA module equivalently into a LoRA module, thereby eliminating the need to modify the original model’s parameters during sharing and usage.
+In the initialization phase, PiSSA decomposes the original matrix into principal components and a residual matrix: $W=W^{res}+A_{0}\times B_{0}$.
+Upon completion of training, the model adjusts the weights as follows: $W+\Delta W=W^{res}+A\times B$.
+Thus, the modification of the model weights by PiSSA is given by:
+$\Delta W$ $=A\times B-A_{0}\times B_{0}$ (6)
+$=[A\,|\,A_{0}]\times\begin{bmatrix}B\\ -B_{0}\end{bmatrix}$ (7)
+$=A^{{}^{\prime}}B^{{}^{\prime}}.$ (8)
+[…]
+Appendix B Training Loss under Various Ranks
+(a) Variation of loss during the fine-tuning of LLaMA 2-7B with PiSSA throughout the training phase.
+(b) Variation of loss during the fine-tuning of Mistral-7B-v0.1 with PiSSA throughout the training phase.
+(c) Variation of loss during the fine-tuning of Gemma-7B with PiSSA throughout the training phase.
+Figure 5: The change in loss during the training process with respect to ranks [2, 4, 8, 16, 32, 64, 128].
+Appendix C Quantization Error for More Layers
+Figure 6: Comparison of quantization errors in QLoRA, LoftQ, and PiSSA across k_proj, v_proj, o_proj and gate_proj, up_proj, down_proj layers.

### v3

@@
 Abstract
-As the parameters of large language models (LLMs) expand, the computational cost of fine-tuning the entire model becomes prohibitive.
-To address this challenge, we introduce a parameter-efficient fine-tuning (PEFT) method, Principal Singular values and Singular vectors Adaptation (PiSSA), which optimizes a significantly reduced parameter space while achieving or surpassing the performance of full-parameter fine-tuning.
-PiSSA is inspired by Intrinsic SAID, which suggests that pre-trained, over-parametrized models inhabit a space of low intrinsic dimension.
-Consequently, PiSSA represents a matrix $W\in\mathbb{R}^{m\times n}$ within the model by the product of two trainable matrices $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, where $r\ll\min(m,n)$, plus a residual matrix $W^{res}\in\mathbb{R}^{m\times n}$ for error correction.
-Singular value decomposition (SVD) is employed to factorize $W$, and the principal singular values and vectors of $W$ are utilized to initialize $A$ and $B$.
-The residual singular values and vectors initialize the residual matrix $W^{res}$, which keeps frozen during fine-tuning.
-Notably, PiSSA shares the same architecture with Low-Rank Adaptation (LoRA), which hypothesizes that changes in model parameters $\Delta W$ form a low-rank matrix.
-However, LoRA approximates $\Delta W$ through the product of two matrices, $A$, initialized with Gaussian noise, and $B$, initialized with zeros, while PiSSA initializes $A$ and $B$ with principal singular values and singular vectors of the original matrix $W$.
-Given that the principal singular values and vectors capture the essence of a low-rank matrix, PiSSA can better approximate the outcomes of full-parameter fine-tuning at the beginning by changing the essential parts while freezing the “noisy” parts.
-In comparison, LoRA freezes the original matrix and updates the “noise”.
-This distinction enables PiSSA to convergence much faster than LoRA and also achieve better performance in the end.
-On five common benchmarks, PiSSA outperforms LoRA on all of them using exactly the same setups except for a different initialization.
-On GSM8K, Mistral-7B fine-tuned with PiSSA achieves an accuracy of 72.86%, outperforming LoRA’s 67.7% by 5.16%.
-Due to the same architecture, PiSSA inherits many of LoRA’s advantages, such as parameter efficiency and compatibility with quantization.
-Furthermore, PiSSA reduces the 4-bit quantization error in LLaMA 2-7B by 18.97%, resulting in a substantial improvement in fine-tuning performance.
-On the GSM8K benchmark, PiSSA achieves an accuracy of 49.13%, surpassing the performances of QLoRA at 39.8% and LoftQ at 40.71%.
-Leveraging a fast SVD technique, the initialization of PiSSA takes only a few seconds, inducing negligible cost of switching LoRA to PiSSA.
+To parameter-efficiently fine-tune (PEFT) large language models (LLMs), the low-rank adaptation (LoRA) method approximates the model changes $\Delta W\in\mathbb{R}^{m\times n}$ through the product of two matrices $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, where $r\ll\min(m,n)$, $A$ is initialized with Gaussian noise, and $B$ with zeros.
+LoRA freezes the original model $W$ and updates the “Noise & Zero” adapter, which may lead to slow convergence.
+To overcome this limitation, we introduce Principal Singular values and Singular vectors Adaptation (PiSSA).
+PiSSA shares the same architecture as LoRA, but initializes the adaptor matrices $A$ and $B$ with the principal components of the original matrix $W$, and put the remaining components into a residual matrix $W^{res}\in\mathbb{R}^{m\times n}$ which is frozen during fine-tuning.
+Compared to LoRA, PiSSA updates the principal components while freezing the “residual” parts, allowing faster convergence and enhanced performance.
+Comparative experiments of PiSSA and LoRA across 12 different models, ranging from 184M to 70B, encompassing 5 NLG and 8 NLU tasks, reveal that PiSSA consistently outperforms LoRA under identical experimental setups.
+On the GSM8K benchmark, Mistral-7B fine-tuned with PiSSA achieves an accuracy of 72.86%, surpassing LoRA’s 67.7% by 5.16%.
+Due to the same architecture, PiSSA is also compatible with quantization to further reduce the memory requirement of fine-tuning.
+Compared to QLoRA, QPiSSA (PiSSA with 4-bit quantization) exhibits smaller quantization errors in the initial stages.
+Fine-tuning LLaMA-3-70B on GSM8K, QPiSSA attains an accuracy of 86.05%, exceeding the performances of QLoRA at 81.73%.
+Leveraging a fast SVD technique, PiSSA can be initialized in only a few seconds, presenting a negligible cost for transitioning from LoRA to PiSSA.
+(a) Full Fine-tuning
+(b) LoRA
+(c) PiSSA
+Figure 1: The comparison among Full Fine-tuning, training with LoRA, and PiSSA.
+In this visualization, blue modules represent parts of the model where parameters are frozen during training, while orange modules indicate components that require updates.
+QLoRA quantizes the pretrained matrix in LoRA to 4-bit, whereas QPiSSA quantizes the residual matrix in PiSSA.
+(a) PiSSA converges more rapidly.
+(b) PiSSA reduces more quantization error.
+Figure 2: We illustrate the two key advantages of PiSSA: converging faster and better, and reducing quantization error.
+In the left figure, we use a toy example to show PiSSA’s faster convergence, where we first train a two-layer MLP classifying odd numbers of MNIST, and then fine-tune the model on even numbers.
+PiSSA finds the right direction more quickly and achieves a lower loss with the same number of steps.
+In the right figure, we show the quantization error reduction of PiSSA compared to directly quantizing the base model.
+PiSSA can reduce more quantization error than the baselines.
 1 Introduction
@@
 As depicted in Figure 1(b), for a pre-trained weight matrix $W\in\mathbb{R}^{m\times n}$, LoRA substitutes the updates with a low-rank decomposition $\Delta W=AB$, where $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, and the rank $r\ll\text{min}(m,n)$.
-For $h=Wx$, the modified forward pass computation is expressed as follows:
-$h=(W+\Delta W)x=(W+AB)x.$ (1)
+For $Y=XW$, the modified forward pass is as follows:
+$Y=X(W+\Delta W)=X(W+AB),$ (1)
+where $X\in\mathbb{R}^{b\times m}$, $Y\in\mathbb{R}^{b\times n}$, and $b$ represents the batch size of input data.
 A random Gaussian initialization is used for $A$ and zero for $B$, making $AB=0$ at the beginning of training, thereby the injection of adapters does not affect the model’s output initially.
-LoRA avoids the need to compute gradients or maintain the optimizer states for the original matrix, instead optimizing the injected, significantly smaller low-rank matrices.
-Thus, it could reduce the number of trainable parameters by 10,000 times and the GPU memory requirement by 3 times.
+LoRA avoids the need to compute gradients or maintain the optimizer states for the original matrix $W$, instead optimizing the injected, significantly smaller low-rank matrices $A,B$.
+Thus, it could reduce the number of trainable parameters by 10,000 times and the GPU memory requirement by 3 times [11].
 Moreover, LoRA often achieves comparable or superior performance to full parameter fine-tuning, indicating that fine-tuning “parts” of the full parameters can be enough for downstream tasks.
-By integrating the quantization of pre-trained models, LoRA also enables reducing the average memory requirements by 16 times [10].
-Meanwhile, the adapters are still allowed to utilize higher precision weights; thus, the quantization usually does not degrade the performance of full-precision base models.
-(a) Full Fine-tuning
-(b) LoRA
-(c) PiSSA
-Figure 1: The comparison among Full Fine-tuning, training with LoRA, and PiSSA.
-In this visualization, blue modules represent parts of the model where parameters are frozen during training, while orange modules indicate components that require updates.
-Inspired by the core idea of Intrinsic SAID [14, 15], our Principal Singular values and Singular vectors Adapter (PiSSA) diverges from LoRA and its successors by not focusing on approximating $\Delta W$, but $W$.
-In alignment with Intrinsic SAID, we assume that the matrix $W$ in pre-trained models exhibits low (intrinsic) rank characteristics.
-Thus, we apply singular value decomposition (SVD) to matrix $W$.
-Based on the magnitude of the singular values, we divide $W$ into two components: the principal matrix $W^{pri}$ and the residual matrix $W^{res}$.
-The principal matrix $W^{pri}$ is of low rank and can be represented by the product of $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, where $r\ll\min(m,n)$.
-$A$ and $B$ are initialized based on the principal singular values and vectors and are trainable, whereas $W^{res}$ is initialized with the product of the residual singular values and vectors and are frozen.
-As indicated by Equation (2), the integration of $AB$ (the PiSSA adaptor) with the residual model also preserves the capability of the pre-trained model in the beginning of fine-tuning:
-$h=Wx=(W^{res}+W^{pri})x=(W^{res}+AB)x.$ (2)
-Different from LoRA, a significantly less important part $W^{res}$ is kept frozen, while the essential part $W^{pri}$ is fully tunable from the beginning, which enables PiSSA to fit the training data faster and better.
-Since the high-frequency principal component $W^{pri}$ is preserved in the adapter, when fine-tuning PiSSA incorporated with quantization, PiSSA significantly reduced quantization errors in the initial stages compared to using QLoRA.
+By integrating the quantization of pre-trained matrices $W$, LoRA also enables reducing the average memory requirements by 16 times [10].
+Meanwhile, the adapters are still allowed to utilize higher precision weights; thus, the quantization usually does not significantly degrade the performance of LoRA.
+According to Equation 1, the gradients of A and B are $\frac{\partial L}{\partial A}=X^{T}\left(\frac{\partial L}{\partial Y}\right)B^{T}$ and $\frac{\partial L}{\partial B}=A^{T}X^{T}\left(\frac{\partial L}{\partial Y}\right)$.
+Compared to full fine-tuning, using LoRA initially does not change the output $Y$ for the same input $X$, so the gradient magnitude is primarily determined by the values of $A$ and $B$.
+Since $A$ and $B$ are initialized with Gaussian noise and zeros in LoRA, the gradients can be very small, leading to slow convergence in the fine-tuning process.
+We also observe this phenomenon empirically, where LoRA often wastes much time around the initial point.
+Our Principal Singular values and Singular vectors Adapter (PiSSA) diverges from LoRA and its successors by focusing not on approximating $\Delta W$, but $W$.
+We apply singular value decomposition (SVD) to matrix $W$.
+Based on the magnitude of the singular values, we partition $W$ into two parts: the principal low-rank matrix $W^{pri}$, comprising a few largest singular values, and the residual matrix $W^{res}$, which possesses the remaining smaller singular values (with a larger quantity, representing a possible long-tail distribution).
+The principal matrix $W^{pri}$ can be represented by the product of $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, where $r\ll\min(m,n)$.
+As depicted in Figure 1(c), $A$ and $B$ are initialized based on the principal singular values and singular vectors and are trainable.
+Conversely, $W^{res}$ is initialized with the product of the residual singular values and singular vectors and remains frozen during fine-tuning.
+Since the principal singular vectors represent the directions in which the matrix $W$ has the most significant stretching or impact, by directly tuning these principal components, PiSSA is able to fit the training data faster and better (as demonstrated in Figure 2(a)).
+Moreover, the loss and gradient norm curves of PiSSA often demonstrate a similar trend to those of full parameter fine-tuning in our experiments (Figure 4), indicating that fine-tuning the principal components matches the behavior of fine-tuning the full matrix to some degree.
+Since the principal components $W^{pri}$ are preserved in the adapter and are of full precision, another benefit of PiSSA is that when applying quantization to the frozen part $W^{res}$, we can significantly reduce the quantization error compared to QLoRA (which quantizes the whole $W$), as illustrated in Figure 2(b).
+Therefore, PiSSA is perfectly compatible with quantization, making it a plug-and-play substitution for LoRA.
 2 Related Works
@@
 Parameter Efficient Fine-Tuning (PEFT) [12, 13] emerges as a compelling solution by minimizing the fine-tuning parameters and memory requirements while achieving comparable performance to full fine-tuning.
-PEFT encompasses strategies like partial fine-tuning, soft prompt fine-tuning, non-linear adapter fine-tuning, and linear adapter fine-tuning.
-Partial fine-tuning selects certain model parts, e.g., bias and layer normalization [16, 17], or inserted trainable mask to the pre-trained weights [18, 19] and delta weights [20, 21, 22, 23], for fine-tuning.
-However, the performance of partial fine-tuning may be limited due to only specific layers’ parameters being fine-tuned.
-Soft prompt fine-tuning  [24, 25, 26, 27, 28, 29, 30] introduces optimizable vectors into the model’s input or hidden layers, adapting to specific tasks with minimal structural changes.
-Non-linear adapter fine-tuning  [31, 32, 33, 34, 35, 36, 37, 38, 39, 40] incorporates (in a non-linear manner) trainable (non-linear) modules to the models.
-Both soft prompt fine-tuning and non-linear adapter fine-tuning add complexity and inference costs despite their adaptability.
-Linear adapter fine-tuning [14, 15, 11, 41], particularly LoRA [11] and its derivative methods, involves adding trainable parameters to the self-attention and feed-forward layers, which are then linearly combined with the original weights.
-After fine-tuning, these adaptations can be reparametrized into the standard model structure, thus gaining widespread adoption due to their ability to maintain the model’s original architecture while enabling efficient fine-tuning.
-LoRA [11] posits that weight adjustments during model adaptation exhibit a low "intrinsic rank".
-It learns the weight updates through a low-rank decomposition, keeping the original structure frozen and proving effective in parameter-efficient fine-tuning.
-Following LoRA, Low-rank Adjustment [42, 41, 43] learns the rank size needed for LoRA in each layer of the model.
-LoRA-guided Pretrained Weight Update [44, 45] updates the original weights of the model using the parameters from adapter layers, enhancing LoRA’s representational capacity.
-Quantization Adaptation [10, 46, 47] applies quantization to the base model, further reducing training consumption overhead.
-Differing from LoRA’s approach of learning low-rank approximations of weight updates, our PiSSA method innovates by utilizing the low intrinsic dimension of pre-trained weight directly.
-It initializes updatable adapter layers with principal singular values and vectors from the model’s weights, preserving the essential capabilities while enabling effective fine-tuning.
-This approach, while different from LoRA’s in philosophy, shares most of the structural benefits of LoRA and can be combined with methods like quantization the residual model for reducing computation cost.
-3 Method
+PEFT encompasses strategies like partial fine-tuning [14, 15, 16, 17, 18, 19, 20, 21], soft prompt fine-tuning  [22, 23, 24, 25, 26, 27, 28], non-linear adapter fine-tuning [29, 30, 31, 32, 33, 34, 35, 36, 37, 38], and low rank adapter based fine-tuning [39, 40, 11, 41].
+LoRA [11] inject trainable adapters to the linear layers.
+After fine-tuning, these adaptations can be re-parameterized into the standard model structure, thus gaining widespread adoption due to their ability to maintain the model’s original architecture while enabling efficient fine-tuning.
+Following LoRA, AdaLoRA [42, 41, 43] dynamically learns the rank size needed for LoRA in each layer of the model.
+DeltaLoRA [44, 45] updates the original weights of the model using parameters from adapter layers, enhancing LoRA’s representational capacity.
+LoSparse [46] incorporates LoRA to prevent pruning from eliminating too many expressive neurons.
+DoRA [47] introduces a magnitude component to learn the scale of $\Delta W$ while utilizing the original AB as a direction component of $\Delta W$.
+Unlike LoRA and its successors, which focus on learning low-rank approximations of weight updates, our PiSSA approach directly tunes the essential but low-rank parts of the model while keeping the noisier, high-rank, and nonessential parts frozen.
+Although our approach differs in philosophy from LoRA, it shares most of LoRA’s structural benefits and can be extended by these methods to enhance its performance.
+QLoRA [10] integrates LoRA with 4-bit NormalFloat (NF4) quantization, along with Double Quantization and Paged Optimizers, enabling the fine-tuning of a 65B parameter model on a single 48GB GPU while preserving the performance of full 16-bit fine-tuning tasks.
+QA-LoRA [48] introduces group-wise operators to increase the degree of freedom in low-bit quantization.
+LoftQ [49] reduces quantization error by decomposing the quantization error matrix of QLoRA and retaining the principal components with an adapter.
+Our PiSSA approach can also be combined with quantization techniques, and we have found that PiSSA significantly reduces quantization error compared to QLoRA and LoftQ.
+3 PiSSA: Principal Singular Values and Singular Vectors Adaptation
 […]
-$A$ $=U_{[:r]}\,\text{diag}(S_{[:r]}^{1/2})\in\mathbb{R}^{m\times r},$ (3)
-$B$ $=\text{diag}(S_{[:r]}^{1/2})\,V_{[:r]}^{T}\in\mathbb{R}^{r\times n}.$ (4)
-The residual singular values and vectors are used to build the residual matrix:
-$W^{res}=U_{[r:]}\,\text{diag}(S_{[r:]})\,V_{[r:]}^{T}\in\mathbb{R}^{m\times n}.$ (5)
-Given $W=W^{res}+AB$, the introduction of PiSSA in the beginning of training does not compromise the model’s capabilities.
-Considering $S_{[:r]}\gg S_{[r:]}$, thus rendering $W^{pri}\gg W^{res}$ and $W\approx AB$, the adapter $AB$ encapsulates the pre-trained model’s primary capabilities.
-Then, the fine-tuning of $AB$ mirrors the process of fine-tuning the entire model, despite using much less parameters.
-In contrast, LoRA initializes the adapter with Gaussian noise and zeros, thus the principal part of $W$ is not updatable in the early stage of the fine-tuning, possibly introducing much waste of gradient descent steps.
+$A$ $=U_{[:,:r]}\,S_{[:r,:r]}^{1/2}\in\mathbb{R}^{m\times r},$ (2)
+$B$ $=S_{[:r,:r]}^{1/2}\,V_{[:,:r]}^{T}\in\mathbb{R}^{r\times n}.$ (3)
+The residual singular values and vectors are used to build the residual matrix which is frozen during fine-tuning:
+$W^{res}=U_{[:,r:]}\,S_{[r:,r:]}\,V_{[:,r:]}^{T}\in\mathbb{R}^{m\times n}.$ (4)
+As indicated by Equation 5, the integration of $AB$ with the residual matrix also preserves the full capability of the pre-trained model in the beginning of fine-tuning:
+$Y=XW=X(W^{res}+W^{pri})=X(W^{res}+AB).$ (5)
+Similar to LoRA, the gradients of $A$ and $B$ are also given by $\frac{\partial L}{\partial A}=X^{T}\left(\frac{\partial L}{\partial Y}\right)B^{T}$ and $\frac{\partial L}{\partial B}=A^{T}X^{T}\left(\frac{\partial L}{\partial Y}\right)$.
+Since elements of $\mathbf{s}_{[:r]}$ $\gg$ elements of $\mathbf{s}_{[r:]}$, the trainable adapter $W^{pri}=AB$ contains the most essential directions of $W$.
+In the ideal case, training $AB$ mirrors the process of fine-tuning the entire model despite using fewer parameters.
+The ability to directly fine-tune the most essential part of a model enables PiSSA to converge faster and better.
+In contrast, LoRA initializes the adapters $A$ and $B$ with Gaussian noise and zeros while keeping $W$ frozen.
+Consequently, the gradients are small or in random directions during the early stages of fine-tuning, possibly introducing much waste of gradient descent steps.
 Moreover, an inferior initialization might lead to suboptimal local minimum points found, causing worse generalization performance.
 Since PiSSA shares the identical architecture with LoRA, it inherits most of LoRA’s benefits.
-These include, but are not limited to, the capability of fine-tuning a model with a reduced number of trainable parameters, quantizing the residual model to decrease memory consumption during forward propagation in training, and easy deployment, during which process PiSSA can significantly reduce the initial quantization error in the initial stages.
+These include but are not limited to the capability of fine-tuning a model with a reduced number of trainable parameters, quantizing the residual model to decrease memory consumption during forward propagation in training, and easy deployment.
 The adapter’s straightforward linear structure facilitates the integration of trainable matrices with the pre-trained weights upon deployment, thereby maintaining the original inference speed of a fully fine-tuned model.
-For storage efficiency, we can choose not to store the dense parameter matrix $\Delta W$, but to store the low-rank matrices, $\Delta AB$ instead.
-As shown in Appendix A, leveraging solely the $\Delta AB$ facilitates their seamless integration with the original pre-trained models.
-Finally, one pre-trained model can accommodate multiple $\Delta AB$, fine-tuned by diverse PiSSA procedures, which enables fast adapting of the pre-trained model to different downstream applications.
+Employing the Fast SVD technique [50] allowed PiSSA to finish initialization in several seconds (Appendix B), which is a negligible cost.
+For storage efficiency, we can choose not to store the dense parameter matrix $\Delta W$, but to store the low-rank matrices, $\Delta A$ and $\Delta B$ instead.
+As shown in Appendix C, leveraging solely the $\Delta A$ and $\Delta B$ facilitates their seamless integration with the original pre-trained models.
+Finally, one pre-trained model can accommodate multiple $\Delta A,\Delta B$, fine-tuned by diverse PiSSA or LoRA procedures, which enables fast adaptation of the pre-trained model to different downstream applications.
+4 QPiSSA: PiSSA with Quantization
+Quantization divides the value range of a matrix into several continuous regions, and maps all values falling inside a region into the same “quantized” value.
+It is an effective technique to reduce the memory consumption of forward propagation, but breaks down during backpropagation.
+At the same time, LoRA greatly reduces the backward memory requirement, making it highly suitable to use LoRA and quantization together, where the base model is quantized for memory-efficient forward propagation, and the LoRA adaptors are kept in full precision for accurate backward parameter updates.
+One representative previous work, QLoRA, quantizes the base model to Normal Float 4-bit (NF4) and initializes the full-precision $A$ and $B$ with Gaussian-Zero initialization.
+Therefore, the overall error is given by:
+Quantization Error of $\text{Quantization Error of QLoRA}=||W-\left(nf4(W)+AB\right)||_{*}=||W-nf4(W)||_{*},$ (6)
+where $||M||_{*}$ denotes the nuclear norm (also known as the trace norm)  [51], defined as:
+$\|M\|_{*}=\operatorname{trace}\left({\sqrt{M^{*}M}}\right)=\sum_{i=1}^{\min\{m,n\}}\sigma_{i}(M),$ (7)
+where $\sigma_{i}(M)$ is the $i^{\text{th}}$ singular value of $M$.
+As we can see, the quantization error of QLoRA is the same as that of directly quantizing the base model.
+Our QPiSSA, however, does not quantize the base model but the residual model.
+Therefore, its error is given by:
+Quantization Error of $\text{Quantization Error of QPiSSA}=||W-\left(nf4(W^{res})+AB\right)||_{*}=||W^{res}-nf4(W^{res})||_{*}.$ (8)
+Since the residual model has removed the large-singular-value components, $W^{res}$ has a narrower distribution than that of $W$, as can be seen in Figures 3(a) and 3(b) (comparing the singular value distributions of $W$ and $W^{res}$), as well as Figures 3(c) and 3(f) (comparing the value distributions of $W$ and $W^{res}$), which is beneficial for reducing the quantization error.
+Furthermore, since NF4 is optimized for normally distributed data [10], we fit a Gaussian for the values of $W$ and $W^{res}$, respectively.
+As can be seen in Figures 3(c) and 3(f), $W^{res}$ is much more Gaussian-like and has a smaller standard deviation, making it more suitable to apply NF4 to $W^{res}$ instead of $W$.
+Both the above lead QPiSSA to achieve a significantly lower quantization error than QLoRA, shown in Figures 3(d) and 3(e).
+(a) Original matrix $W$
+(b) Residual matrix $W^{res}$
+(c) The distribution of $W$
+(d) Error matrix of QLoRA
+(e) Error matrix of QPiSSA
+(f) The distribution of $W^{res}$
+Figure 3: Figures (a), (b), (d) and (e) visualize the singular values in descending order for $W$, $W^{res}$, $W-nf4(W)$ and $W^{res}-nf4(W^{res})$ of LLaMA 2-7B’s “layers[0].self_attn.q_proj” matrix.
+Figures (c) and (f) show the data value distributions of $W$ and $W^{res}$.
+Besides the advantage of reducing quantization error, QPiSSA’s gradient direction is similar to that of PiSSA, resulting in significantly better fine-tuning performance compared to QLoRA.
 5 Experiments
-We conducted comparative experiments among PiSSA, LoRA, and full parameter fine-tuning on three models: LLaMA 2-7B [48], Mistral-7B-v0.1 [49], and Gemma-7B [50].
-Our comparison spanned a diverse range of tasks, encompassing mathematics [2, 51, 52], coding [4, 53, 54, 55], and instruction tuning [5, 7, 6].
-PiSSA exhibited superior performance across these downstream tasks (Section 4.1).
-When combined with quantization techniques, PiSSA effectively reduces quantization errors.
-Compared to QLoRA [10] and LoftQ [47], PiSSA exhibits superior performance both during initialization and fine-tuning stages (Section 4.2).
-The ablation studies, which varied across different ranks, showed that PiSSA converged more swiftly and aligned more closely with the training data than LoRA, demonstrating robust superiority under equivalent trainable parameter configurations (Section 4.3).
-Further experiments demonstrated that employing the Fast SVD technique [56] allowed PiSSA to achieve a balance between initialization speed and performance (Section 4.4).
-Lastly, adapters initialized with principal singular values and vectors (PiSSA) were compared to those initialized with intermediate or less significant ones, revealing that direct fine-tuning of the model’s principal components yielded better outcomes (Section 4.5).
+The experiments were conducted on the NVIDIA A800-SXM4(80G) GPU.
 In our experiments, we adopt the Alpaca [52] implementation strategy, using the AdamW optimizer with a batch size of 128, a learning rate of 2e-5, cosine annealing schedules, and a warmup ratio of 0.03, without any weight decay.
-We utilize the bf16 computation type, ensure lora_alpha is always equal to lora_r, set lora_dropout to 0, and incorporate the adapters into all linear layers of the base model.
-We train only on the first 100,000 conversions from the dataset for one epoch, and only using the response of the conversions to compute loss.
-The experiments in this section were conducted on the NVIDIA A800-SXM4(80G) GPU.
-4.1 Comparative Analysis of PiSSA, LoRA, and Full Fine-Tuning Across Various Tasks
-Table 1 presents a comparative evaluation of fine-tuning strategies using PiSSA, LoRA, and full parameter fine-tuning.
-In these experiments, we fine-tune LLaMA 2-7B, Mistral-7B-v0.1, and Gemma-7B on the MetaMathQA dataset [2] to assess their mathematical problem-solving capabilities on the GSM8K [51] and MATH [2] validation sets.
-Additionally, the models are fine-tuned on the CodeFeedback dataset [53] and their coding proficiency is evaluated on the HumanEval [54] and MBPP [55] datasets.
-Furthermore, the models are trained on the WizardLM-Evol-Instruct dataset [7] and tested for their conversational abilities on the MT-Bench dataset [6].
-Table 1: Performance of different models on various tasks
+As discussed in Section B.3 of QLoRA [10], we compute the loss using only the responses from the instruction-following datasets.
+We ensure lora_alpha is always equal to lora_r, set lora_dropout to 0, and incorporate the adapters into all linear layers of the base model.
+We utilize the Float32 computation type for both the base model and the adapter in LoRA and PiSSA.
+For QLoRA, LoftQ, and QPiSSA, we use 4-bit NormalFloat [10] for the base model and Float32 for the adapter.
+BFloat16 [53] is used for full parameter fine-tuning to save the resources (see Appendix D).
+5.1 Evaluating the Performance of PiSSA on both NLG and NLU Tasks
+Table 1 presents a comparative evaluation of fine-tuning strategies PiSSA, LoRA, and full parameter fine-tuning on natural language generation (NLG) tasks.
+We fine-tuned LLaMA 2-7B [54], Mistral-7B-v0.1 [55], and Gemma-7B [56] on the MetaMathQA dataset [2] to assess their mathematical problem-solving capabilities on the GSM8K [57] and MATH [2] validation sets.
+Additionally, the models were fine-tuned on the CodeFeedback dataset [58] and evaluated for coding proficiency using the HumanEval [59] and MBPP [60] datasets.
+Furthermore, the models were trained on the WizardLM-Evol-Instruct dataset [7] and tested for conversational abilities on the MT-Bench dataset [6].
+All experiments were conducted using subsets containing 100K data points and were trained for only one epoch to reduce training overhead.
+Table 1: Comparing PiSSA and LoRA on NLG tasks
 Model Strategy Trainable GSM8K MATH HumanEval MBPP MT-Bench
 Parameters
-LLaMA 2-7B Full Fine-tuning 6738M 49.05 7.22 21.34 35.59 4.91
-LoRA 320M 42.3 5.5 18.29 35.34 4.58
+LLaMA 2-7B Full FT 6738M 49.05 7.22 21.34 35.59 4.91
+LoRA 320M 42.30 5.50 18.29 35.34 4.58
 PiSSA 320M 53.07 7.44 21.95 37.09 4.87
-Mistral-7B Full Fine-tuning 7242M 67.02 18.6 45.12 51.38 4.95
-LoRA 168M 67.7 19.68 43.9 58.39 4.9
+Mistral-7B Full FT 7242M 67.02 18.6 45.12 51.38 4.95
+LoRA 168M 67.70 19.68 43.90 58.39 4.90
 PiSSA 168M 72.86 21.54 46.95 62.66 5.34
-Gemma-7B Full Fine-tuning 8538M 71.34 22.74 46.95 55.64 5.4
-LoRA 200M 74.9 31.28 53.66 65.41 4.98
+Gemma-7B Full FT 8538M 71.34 22.74 46.95 55.64 5.40
+LoRA 200M 74.90 31.28 53.66 65.41 4.98
 PiSSA 200M 77.94 31.94 54.27 66.17 5.64
-As can be seen from Table 1, across all models and tasks, fine-tuning with PiSSA surpasses the effects of fine-tuning with LoRA, as well as most of full fine-tuning.
-In the future, we will extend the comparing to more models and tasks.
-4.2 Comparative Analysis of QLoRA, LoftQ, and PiSSA in Quantization Fine-tuning
-In this section, we compare the performance of QLoRA, LoftQ, and PiSSA in fine-tuning with quantization.
-All base models are quantized to ‘nf4’ precision and frozen, while the adapters are trained using ‘bfloat16’ precision.
-We begin by comparing the quantization errors of the three methods.
-For each method, the quantized baseline model is combined with its adapter, and the difference between the original weights and this combination serves as the error matrix.
-We measure the magnitude of the errors by the average nuclear norm of all error matrices.
-Using the quantization error of the original model without any adapter as a baseline, the errors for QLoRA, LoftQ, and PiSSA are divided by this baseline to express them as error rates.
-As QLoRA’s adapter employs Gaussian-zero initialization, the error rate does not decrease; LoftQ can eliminate quantization errors through multiple singular value decompositions, but for consistency with PiSSA, only one decomposition is performed.
-PiSSA requires a full-precision singular value decomposition, storing the principal singular values and vectors in the adapter before quantizing the residual model.
-We train LLaMA 2-7B on MetaMathQA and validate the effectiveness of the three methods on GSM8K and MATH datasets.
-The reductions in error and the training results are displayed in Figure 2.
-Figure 2: Comparing the quantization error, the fine-tuning loss on the MetaMathQA and the accuracy on the GSM8K and MATH validation sets.
-The results, shown in Figure 2, reveal that PiSSA significantly outperforms LoftQ in reducing errors, achieving a 19% reduction at rank 128, as opposed to LoftQ’s 8%.
-LoftQ’s fine-tuning improvements over QLoRA are not significant, and in certain settings, it even underperforms compared to QLoRA.
-PiSSA effectively reduces quantization errors, and the training of principal singular values and vectors additionally provides benefits, thus substantially enhancing training outcomes in comparison to QLoRA and LoftQ.
-4.3 Evaluating PiSSA and LoRA Performance Across Diverse Ranks
-This section explores the impact of incrementally increasing the rank of PiSSA and LoRA from 1 to 128, aiming to determine the optimal number of trainable parameters needed to achieve performance comparable to full parameter training.
-The training is conducted using the MetaMathQA dataset, while the validation is performed on the GSM8K and MATH datasets.
-The outcomes of these experiments are depicted in Figure 3, with additional results presented in Figure 5 in the Appendix.
-(a) Variation of loss with respect to rank 1 throughout the training phase. Additional ranks are depicted in Figure 5 in the Appendix.
-(b) Final training loss across different ranks.
-(c) Rank-wise performance evaluated using pass@1 on the GSM8K dataset.
-(d) Rank-wise performance evaluated using pass@1 on the MATH dataset.
-Figure 3: Fine-tuning LLaMA 2-7B, Mistral-7B-v0.1, and Gemma-7B on the MetaMathQA dataset: A comparison of full parameter fine-tuning (indicated by a dashed line), LoRA (in blue), and PiSSA (in orange).
-From Figure 3(a), it is observed that at the beginning of training, PiSSA quickly starts to converge, whereas LoRA experiences a period where the loss not only fails to decrease but even slightly increases.
-Furthermore, throughout the entire training process, the loss of PiSSA consistently remains at a lower level compared to LoRA.
-Figure 3(b) shows the final loss on the training set for models trained with ranks varying from 1 to 128.
-The data indicate that, compared to LoRA, PiSSA achieves a level of fitting to the training set similar to that of full-parameter fine-tuning with fewer trainable parameters.
-In Figures 3(c) and 3(d), we compare the performance of fine-tuned models on the GSM8K and MATH validation sets under various ranks, finding that PiSSA outperforms LoRA with the same amount of trainable parameters.
-Furthermore, as the rank increases, PiSSA quickly reaches and surpasses the performance of full-parameter fine-tuning.
-Notably, when the training loss of LLaMA 2-7B with a rank of 128 PiSSA equals that of full-parameter fine-tuning, it surpasses full-parameter fine-tuning in terms of generalization on downstream tasks.
-This validates our approach of using principal singular values and vectors to replace full-parameter model fine-tuning, ignoring components of lower rank, resulting in a model with improved robustness.
-4.4 Fast Singular Value Decomposition
+As shown in Table 1, across all models and tasks, fine-tuning with PiSSA consistently surpasses the performance of fine-tuning with LoRA.
+For instance, fine-tuning the LLaMA, Mistral, and Gemma with PiSSA results in performance improvements of 10.77%, 5.26%, and 3.04% respectively in mathematical tasks.
+In coding tasks, the improvements are 20%, 6.95%, and 1.14% respectively.
+In MT-Bench, we observe enhancements of 6.33%, 8.98%, and 13.25%.
+Notably, using PiSSA with only 2.3% of Gemma’s trainable parameters outperforms full parameter fine-tuning by 15.59% in coding tasks.
+Further experiments demonstrated that this improvement is robust across various amounts of training data and epochs (Section 5.2), including both 4-bit and full precision (Section 5.3), different model sizes and types (Section 5.4), and varying proportions of trainable parameters (Section 5.5).
+Table 2: Comparing PiSSA and LoRA on NLU tasks. Details can be found in Appendix  I.
+Method Parameters MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B
+RoBERTa-large (355M)
+Full FT 355M 90.2 96.4 90.9 68.0 94.7 92.2 86.6 91.5
+LoRA 1.84M 90.6 96.2 90.9 68.2 94.9 91.6 87.4 92.6
+PiSSA 1.84M 90.7 96.7 91.9 69.0 95.1 91.6 91.0 92.9
+DeBERTa-v3-base (184M)
+Full FT 184M 89.90 95.63 89.46 69.19 94.03 92.40 83.75 91.60
+LoRA 1.33M 90.65 94.95 89.95 69.82 93.87 91.99 85.20 91.60
+PiSSA 1.33M 90.43 95.87 91.67 72.64 94.29 92.26 87.00 91.88
+We also evaluate PiSSA’s natural language understanding (NLU) capability on the GLUE benchmark [61] with RoBERTa-large [62] and DeBERTa-v3-base [63].
+Table 2 presents the results across 8 tasks conducted with the two base models.
+PiSSA outperforms LoRA on 14 out of 16 experimental settings, and achieves the same performance on QQP + RoBERTa.
+The single exception arises from the DeBERTa-based model on the MNLI dataset.
+Upon reviewing the training loss, we observed that PiSSA’s average loss of $0.17$ was lower than LoRA’s $0.24$ in the final epoch.
+This indicates that the fitting ability of PiSSA remains stronger than that of LoRA.
+5.2 Experiments using Full Data and More Epochs
+In this section, we finetune LLaMA 2-7B model on the complete MetaMathQA-395K dataset for 3 epochs to ensure thorough saturation.
+The training loss and gradient norms is visualized to demonstrate quicker convergence and evaluated on the GSM8K dataset every 1000 steps to demonstrate superior performance of PiSSA compared to LoRA.
+The results are depicted in Figure 4.
+Additionally, similar comparisons on Mistral-7B and Gemma-7B are detailed in Appendix G.
+(a) Loss over steps.
+(b) Grad norm over steps.
+(c) Accuracy on GSM8K over training steps.
+Figure 4: The loss, grad norm, and evaluation accuracy over the training steps of LoRA (indicated in blue), PiSSA (in orange), and full parameter fine-tuning (in green).
+According to Figure 4(a), the loss of PiSSA reduces rapidly during the first 100 steps, and the grad norm (shown in Figure 4(b)) of PiSSA is significantly higher than that of LoRA, with a trend similar to full fine-tuning.
+Throughout the process, the loss of PiSSA remains lower than that of LoRA, indicating that PiSSA converges to a better local optimum.
+As shown in Figure 4(c), PiSSA consistently achieves higher accuracy compared to LoRA, and in most cases also surpasses full parameters fine-tuning.
+We hypothesize that this is because PiSSA is a denoised version of full fine-tuning.
+Comparing the grad norm and loss curves of PiSSA and full fine-tuning, we can see that the larger grad norm of full fine-tuning does not bring lower loss, indicating that a portion of the grad norm is spent on noisy directions not beneficial for loss reduction.
+This phenomenon is consistent with Figure 2(a).
+5.3 Conducting 4-bit Quantization Experiments
+In this section, we first compare the initial quantization error reduction ratio of PiSSA, QLoRA, and LoftQ.
+This ratio is defined as $(1-\frac{||W-(nf4(W^{{}^{\prime}})+AB)||_{*}}{||W-nf4(W)||_{*}})\times 100\%$, measuring the relative error decrease achieved by each mehod compared to directly quantizing the base model.
+The partial results are presented in Table 3, and the complete results can be found in Table 6 in Appendix E.
+Table 3: The quantization error reduction ratio of QLoRA, LoftQ, and PiSSA across different layers.
+Method Rank Q K V O Gate Up Down AVG
+LLaMA 2-7B QLoRA 0 0 0 0 0 0 0 0
+loftQ 128 16.5 16.5 15.9 16.0 12.4 12.4 12.3 14.6
+PiSSA 128 27.9 27.2 18.7 18.6 15.8 13.6 13.6 19.4
+LLaMA 3-8B QLoRA 0 0 0 0 0 0 0 0
+loftQ 128 16.4 29.8 28.8 16.1 11.9 11.7 11.7 18.1
+PiSSA 128 26.3 41.7 32.3 20.1 14.4 12.5 12.9 22.9
+LLaMA 3-70B QLoRA 0 0 0 0 0 0 0 0
+LoftQ 64 6.1 17.8 17.0 6.0 4.3 4.4 4.2 8.5
+PiSSA 64 15.7 34.2 18.9 7.5 6.7 5.7 4.7 13.4
+PiSSA 128 23.2 49.0 30.5 12.5 10.1 8.8 8.2 20.3
+In Table 3, PiSSA reduces the quantization error by about 20% compared to directly quantizing the base model.
+The reduction is more significant for lower-rank matrices.
+For instance, in the LLaMA-3-70B [64], all “Key” projection layers see a reduction of 49%.
+The results in Table 3 validate that QLoRA, discussed in Section 4, does not reduce quantization error.
+In contrast, PiSSA significantly outperforms LoftQ in reducing quantization error, as further discussed in Appendix F.
+Besides reducing the quantization error, we expect QPiSSA to also converge faster than QLoRA and LoftQ.
+We train LLaMA 3-8B with LoRA/QLoRA, PiSSA/QPiSSA, LoftQ, and full fine-tuning on MetaMathQA-395K for 3 epochs and record the loss, grad norm, and accuracy on GSM8K.
+(a) Loss over training steps.
+(b) Grad norm over training steps.
+(c) Accuracy on GSM8K over training steps.
+Figure 5: The loss, grad norm, and evaluation accuracy over the training steps of (Q)LoRA, (Q)PiSSA, LoftQ and full parameter fine-tuning.
+According to Figure 5, QPiSSA’s loss reduction speed in the first 100 steps is even faster than PiSSA and full fine-tuning.
+Although LoftQ can reduce the quantization error, its loss convergence speed is not faster than LoRA and QLoRA, indicating that QPiSSA’s ability to reduce the quantization error and its fast convergence might also be orthogonal capabilities.
+After sufficient training, QPiSSA’s loss is also much lower than that of LoRA/QLoRA and LoftQ.
+The grad norm is significantly larger than those of LoRA/QLoRA and LoftQ.
+In terms of fine-tuning performance, QPiSSA’s accuracy is higher than that of QLoRA and LoftQ and even better than that of full-precision LoRA.
+5.4 Experiments Across Various Sizes and Types of Models
+In this section, we compare (Q)PiSSA and (Q)LoRA across 9 models, ranging from 7-70B parameters, including LLaMA 2-7/13B [54], LLaMA-3-8/70B [64], Mistral-7B [55], Gemma-7B [56], and Qwen1.5-7B [65], Yi-1.5-34B [66] and MoE models: DeepSeek-MoE-16B [67] and Mixtral-8x7B [68].
+These models were fine-tuned on the MetaMathQA-100K and CodeFeedback-100K dataset and evaluated on the GSM8K and HumanEval.
+DeepSeek-MoE-16B, Mixtral-8x7B, Yi-1.5-34B, and LLaMA-3-70B were fine-tuned with QPiSSA and QLoRA, while the other models were using PiSSA and LoRA.
+From Figure 6, it can be observed that (Q)PiSSA, compared to (Q)LoRA, shows improved accuracy across various sizes and types of models, demonstrating its consistent advantage over (Q)LoRA.
+Figure 6: Comparison of (Q)PiSSA and (Q)LoRA across models from 7B to 70B.
+5.5 Experiments on Various Ranks
+This section explores the impact of incrementally increasing the rank of PiSSA/QPiSSA and LoRA/QLoRA from 1 to 128, aiming to determine whether PiSSA/QPiSSA consistently outperforms LoRA/QLoRA under different ranks.
+The training is conducted using the MetaMathQA-100K dataset for 1 epoch, while the validation is performed on the GSM8K and MATH datasets.
+The outcomes of these experiments are depicted in Figure 7, with additional results presented in Appendix H.
+(a) Quantization error reduction ratio across ranks.
+(b) Training loss under various ranks.
+(c) Accuracy on GSM8K under various ranks.
+(d) Accuracy on MATH under various ranks.
+Figure 7: The comparison among (Q)LoRA, (Q)PiSSA, LoftQ, and full fine-tuning across ranks.
+Figure 7(a) illustrates the quantization error reduction ratio across various ranks.
+In this figure, QLoRA shows no reduction in quantization error, while QPiSSA consistently outperforms LoftQ in reducing quantization error across all ranks, with a particularly notable advantage at lower ranks.
+In Figure 7(b), the final loss on the training set is shown for models trained with ranks ranging from 1 to 128.
+The results indicate that PiSSA and QPiSSA achieve a better fit to the training data compared to LoRA, QLoRA, and LoftQ.
+In Figures 7(c) and Figures 7(d), we compare the accuracy of the fine-tuned models on the GSM8K and MATH validation sets under various ranks, finding that PiSSA consistently outperforms LoRA with the same amount of trainable parameters.
+Furthermore, as the rank increases, PiSSA will reach and surpass the performance of full-parameter fine-tuning.
+6 Conclusion
+This paper presents a PEFT technique that applies singular value decomposition (SVD) to the weight matrix of pre-trained models.
+The principal components obtained from the SVD are used to initialize a low-rank adapter named PiSSA, while the residual components are kept frozen, in order to achieve effective fine-tuning and parameter efficiency at the same time.
+Through extensive experiments, we found that PiSSA and its 4-bit quantization version QPiSSA significantly outperform LoRA and QLoRA in both NLG and NLU tasks, across different training steps, various model sizes and types, and under various amount of trainable parameters.
+PiSSA provides a novel direction for research in PEFT by identifying and fine-tuning the principal components within the model, analogous to slicing and re-baking the richest slice of a pizza.
+As PiSSA shares the same architecture as LoRA, it can be seamlessly used in existing LoRA pipelines as an efficient alternative initialization method.
+7 Limitation
+There are still some questions with PiSSA not addressed in this paper: 1) Besides language models, can PiSSA also be adapted to convolutional layers and enhance the performance of vision tasks?
+2) Can PiSSA also benefit from some improvements to LoRA, such as AdaLoRA [69] and DyLoRA [42] which adaptively adjust the rank?
+3) Can we provide more theoretical explanations for the advantages of PiSSA over LoRA?
+We are actively exploring these questions.
+Nevertheless, we are excited to see the huge potential of PiSSA already demonstrated in existing experiments and look forward to more tests and suggestions from the community.
+Appendix A Conductive Experiments on Various SVD Components
+To investigate the influence of singular values and vectors of varying magnitudes on the fine-tuning performance, we initialize the adapters injected into LLaMA 2-7B, Mistral-7B-v0.1, and Gemma-7B with principal, medium, and minor singular values and vectors.
+These models are then fine-tuned on the MetaMathQA dataset [2] and evaluated against the GSM8K [57] and MATH datasets [70], with the outcomes depicted in Figures 8.
+Figure 8: Initializing with principal, medium, and minor singular values and vectors, the training loss on the MetaMathQA and the accuracy on the GSM8K and MATH validation sets are reported, respectively, for three models.
+The results highlight that initializing adapters with principal singular values and vectors consistently leads to reduced training loss and enhanced accuracy on both the GSM8K and MATH validation datasets across all three models.
+This underscores the efficacy of our strategy in fine-tuning the model parameters based on the principal singular values.
+Appendix B Fast Singular Value Decomposition
 In order to speed up the decomposition of the pre-trained matrix $W$, we adopted the algorithm proposed by Halko et.al [50] (denoted as Fast SVD), which introduces randomness to achieve an approximate matrix decomposition.
-We compare the initialization time, error, and training loss between SVD and Fast SVD, with the results shown in Table 4.4.
+We compare the initialization time, error, and training loss between SVD and Fast SVD, with the results shown in Table B.
 Initialization time refers to the computation time taken to decompose the pre-trained parameter matrix $W$, measured in seconds.
@@
 1 2 4 8 16 32 64 128
-Initialization Time 1 5.05 8.75 5.07 8.42 5.55 8.47 6.80 11.89
+Initialize Time 1 5.05 8.75 5.07 8.42 5.55 8.47 6.80 11.89
 2 4.38 4.71 4.79 4.84 5.06 5.79 7.70 16.75
@@
 $\infty$ 434.92 434.15 434.30 435.42 435.25 437.22 434.48 435.84
-Initialization Error 1 1.30E-3 1.33E-3 1.55E-3 1.9E-3 1.98E-3 1.97E-3 2.00E-3 1.93E-3
+Initialize Error 1 1.30E-3 1.33E-3 1.55E-3 1.9E-3 1.98E-3 1.97E-3 2.00E-3 1.93E-3
 2 5.84E-4 1.25E-3 1.45E-3 1.43E-3 1.48E-3 1.55E-3 1.48E-3 1.33E-3
@@
 At the same rank, with the increase of niter, the training loss of models initialized based on Fast SVD approaches that of models initialized based on SVD.
-4.5 Verifying the importance of principal singular values and vectors
-To investigate the influence of singular values and vectors of varying magnitudes on the fine-tuning performance, we initialize the adapters injected into LLaMA 2-7B, Mistral-7B-v0.1, and Gemma-7B with principal, medium, and minor singular values and vectors.
-These models are then fine-tuned on the MetaMathQA dataset [2] and evaluated against the GSM8K [51] and MATH datasets [52], with the outcomes depicted in Figures 4.
-Figure 4: Initializing with principal, medium, and minor singular values and vectors, the training loss on the MetaMathQA and the accuracy on the GSM8K and MATH validation sets are reported, respectively, for three models.
-The results highlight that initializing adapters with principal singular values and vectors consistently leads to reduced training loss and enhanced accuracy on both the GSM8K and MATH validation datasets across all three models.
-This underscores the efficacy of our strategy in fine-tuning the model parameters based on the principal singular values.
-5 Conclusion
-This paper introduces a parameter-efficient fine-tuning (PEFT) technique in which singular value decomposition is applied to the weight matrix of pre-trained models to extract principal components.
-These components are then used to initialize an adapter named PiSSA.
-Fine-tuning PiSSA can closely replicate the effects of fine-tuning the complete model in the beginning, while maintaining a good parameter efficiency.
-Through extensive experiments, we show that PiSSA’s fine-tuning performance significantly surpasses that of LoRA.
-PiSSA provides a novel direction for research in PEFT by identifying and fine-tuning the principal components within the model, analogous to slicing and re-baking the richest slice of a pizza.
-As PiSSA shares the same architecture as LoRA, it is easy to use and provides an efficient alternative initialization method.
-There remain several questions to be verified in the future: 1) Can PiSSA achieve significant improvements on a broader range of tasks and larger models?
-2) When the iteration steps of LoRA are sufficiently long (adequately fitting the data), can it match the performance of PiSSA?
-3) Can combining the successors of LoRA with PiSSA lead to further enhancement?
-4) How to give a theoretical explanation for the advantages of PiSSA over LoRA?
-We are actively exploring these questions.
-Nevertheless, we are excited to see the huge potential of PiSSA already demonstrated in existing experiments, and look forward to more tests and suggestions from the community.
-Appendix A Equivalently Convert PiSSA into LoRA
+Appendix C Equivalently Converting PiSSA into LoRA
 The advantage of PiSSA lies in its ability to significantly enhance training outcomes during the fine-tuning phase.
 After training, it allows for the direct sharing of the trained matrices $A$ and $B$.
-However, users need to perform singular value decomposition on the original model, which requires additional time.
-When employing fast singular value decomposition, there can be slight inaccuracies.
-More importantly, utilizing PiSSA for inference necessitates altering the parameters of the original model, which can be inconvenient when using multiple adapters, especially when some adapters might be disabled or activated.
+However, if we directly save $A,B$, users need to perform singular value decomposition on the original model to get $W^{res}$, which requires additional time.
+When employing fast singular value decomposition, there can be slight inaccuracies too.
+More importantly, such a way necessitates altering the parameters of the original model, which can be inconvenient when using multiple adapters, especially when some adapters might be disabled or activated.
 Therefore, we recommend converting the trained PiSSA module equivalently into a LoRA module, thereby eliminating the need to modify the original model’s parameters during sharing and usage.
-In the initialization phase, PiSSA decomposes the original matrix into principal components and a residual matrix: $W=W^{res}+A_{0}\times B_{0}$.
-Upon completion of training, the model adjusts the weights as follows: $W+\Delta W=W^{res}+A\times B$.
+In the initialization phase, PiSSA decomposes the original matrix into principal components and a residual matrix: $W=W^{res}+AB$.
+Upon completion of training, the model adjusts the weights as follows: $W+\Delta W=W^{res}+A^{\prime}B^{\prime}$.
 Thus, the modification of the model weights by PiSSA is given by:
-$\Delta W$ $=A\times B-A_{0}\times B_{0}$ (6)
-$=[A\,|\,A_{0}]\times\begin{bmatrix}B\\ -B_{0}\end{bmatrix}$ (7)
-$=A^{{}^{\prime}}B^{{}^{\prime}}.$ (8)
-[…]
-Appendix B Training Loss under Various Ranks
-(a) Variation of loss during the fine-tuning of LLaMA 2-7B with PiSSA throughout the training phase.
-(b) Variation of loss during the fine-tuning of Mistral-7B-v0.1 with PiSSA throughout the training phase.
-(c) Variation of loss during the fine-tuning of Gemma-7B with PiSSA throughout the training phase.
-Figure 5: The change in loss during the training process with respect to ranks [2, 4, 8, 16, 32, 64, 128].
-Appendix C Quantization Error for More Layers
+$\Delta W$ $=A^{\prime}B^{\prime}-AB$ (9)
+$=\underbrace{[A^{\prime}~{}A]}_{\Delta A}\underbrace{\begin{bmatrix}B^{\prime}\\ -B\end{bmatrix}}_{\Delta B}$ (10)
+where $\Delta A\in\mathbb{R}^{m\times 2r}$ and $\Delta B\in\mathbb{R}^{2r\times n}$.
+Therefore, we can store and share the new adaptor $\Delta A$ and $\Delta B$ instead of $A^{\prime},B^{\prime}$, which allows directly inserting the adaptor to the original matrix and avoids breaking $W$.
+Since $r$ is typically small, the twice storage overhead is still acceptable.
+This modification allows for plug-and-play usage without the need for singular value decomposition, saving time and avoiding computational errors associated with the SVD, without necessitating changes to the original model parameters.
+Appendix D Comparison of Fine-Tuning in BF16 and FP32 Precision
+In this section, we compare the effects of training with BFloat16 and Float32 precision.
+The comparing include four models: LLaMA-2-7B, Mistral-7B, Gemma-7B, and LLaMA-3-8B, each fine-tuned with all parameters in both BFloat16 and Float32 precision on the MetaMathQA-395K dataset.
+The validation results conducted on the GSM8K dataset are shown in Figure 5.
+Table 5: Comparison of fine-tuning results of LLaMA-2-7B, Mistral-7B, Gemma-7B, and LLaMA-3-8B in BF16 and FP32 precision on MetaMathQA-395K dataset for 3 epochs.
+Model Training Loss GSM8K ACC (%) MATH ACC (%)
+BF16 FP32 BF16 FP32 BF16 FP32
+LLaMA-2-7B 0.1532 0.1316 63.15 68.31 13.14 20.38
+Mistral-7B 0.1145 0.1306 73.09 65.88 26.44 23.66
+Gemma-7B 0.1331 0.1382 75.21 75.97 29.18 28.64
+LLaMA-3-8B 0.1271 0.1317 81.96 75.44 33.16 28.72
+From Table 5, it is evident that the choice of precision greatly affects the experimental results.
+For example, the LLaMA-2-7B model shows a 5.16% higher performance on the GSM8K dataset when using FP32 compared to BF16.
+Conversely, the Mistral-7B and LLaMA-3-8B on GSM8K are 7.21% and 6.52% lower with FP32 than with BF16 separately.
+The Gemma-7B model shows similar performance with both precisions.
+Unfortunately, the experiments did not prove which precision is better.
+To reduce training costs, we use BF16 precision when fine-tuning all parameters.
+For methods with lower training costs, such as LoRA, PiSSA, we use FP32 precision.
+For QLoRA, QPiSSA and LoftQ, the base model was used NF4 precision, while the adapter layers used FP32 precision.
+Appendix E Reducing Quantization Error through QPiSSA with Multiple Iteration of SVD
+Table 6 provides a supplementary explanation of the results in Table 3.
+When number of iterations $T>1$, LoftQ uses an $N$-bit quantized weight $Q\in\mathbb{R}_{N}^{m\times n}$ and low-rank approximations $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{n\times r}$ to minimize the following objective by alternating between quantization and singular value decomposition:
+$\underset{Q,A,B}{\min}\|W-(Q+AB^{\top})\|_{F},$ (11)
+where $\|\cdot\|_{F}$ denotes the Frobenius norm, $A$ and $B$ are set to zero. Inspired by LoftQ, our QPiSSA $T$-iter alternately minimize the following objective:
+$\underset{W_{res},A,B}{\min}\|W-(nf4(W_{res})+AB^{\top})\|_{F},$ (12)
+where $A$ and $B$ are initialized by the principal singular values and singular vectors. The process is summarized in Algorithm 1:
+Algorithm 1 QPiSSA-$T$-iters, $T\geq 2$
+0:  Pre-trained weight $W$, target rank $r$, 4-bit quantization function $nf4(\cdot)$, alternating step $T$
+1:  Initialize $A_{0},B_{0}\leftarrow\text{SVD}(W)$ by (2) and (3)
+2:  Initialize residual weight $W_{res}\leftarrow W-A_{0}B_{0}^{\top}$
+3:  for t = $2$ to $T$ do
+4:     Update $A_{t},B_{t}\leftarrow\text{SVD}(W-nf4(W_{res}))$ by (2) and (3)
+5:     Update residual weight $W_{res}\leftarrow W-A_{t-1}B_{t-1}^{\top}$
+6:  end for
+6:  $nf4(W_{res}),A_{T},B_{T}$
+Table 6: PiSSA reduces more quantization error on various ranks and number of iterations.
+Method Rank niter Q K V O Gate Up Down AVG
+LLaMA -2-7B QLoRA 0 0 0 0 0 0 0 0
+loftQ 128 1 8.1 8.1 7.2 7.3 5.3 5.1 5.1 6.6
+PiSSA 128 1 19.0 18.1 8.9 8.9 8.2 5.9 6.0 10.7
+loftQ 128 5 16.5 16.5 15.9 16.0 12.4 12.4 12.3 14.6
+PiSSA 128 5 27.9 27.2 18.7 18.6 15.8 13.6 13.6 19.4
+LLaMA -3-8B QLoRA 0 0 0 0 0 0 0 0
+LoftQ 64 1 4.3 11.0 9.9 3.9 2.7 2.5 2.6 5.3
+PiSSA 64 1 11.3 16.4 8.8 6.3 4.5 2.9 3.3 7.7
+loftQ 64 5 10.1 18.8 18.2 9.9 7.1 7.1 7.1 11.2
+PiSSA 64 5 17.1 27.3 19.5 12.1 8.9 7.2 7.6 14.3
+loftQ 128 1 8.2 20.7 18.8 7.5 5.2 4.8 4.9 10.0
+PiSSA 128 1 17.1 26.5 10.7 10.7 7.0 5.0 5.6 11.8
+loftQ 128 5 16.4 29.8 28.8 16.1 11.9 11.7 11.7 18.1
+PiSSA 128 5 26.3 41.7 32.3 20.1 14.4 12.5 12.9 22.9
+LLaMA -3-70B QLoRA 0 0 0 0 0 0 0 0
+LoftQ 64 1 2.4 11.6 9.2 1.9 1.8 1.7 1.3 4.3
+PiSSA 64 1 12.3 25.0 9.0 4.1 4.2 3.2 2.2 8.6
+LoftQ 64 5 6.1 17.8 17.0 6.0 4.3 4.4 4.2 8.5
+PiSSA 64 5 15.7 34.2 18.9 7.5 6.7 5.7 4.7 13.4
+PiSSA 128 1 17.7 36.6 15.7 6.7 5.8 4.5 3.8 13.0
+PiSSA 128 5 23.2 49.0 30.5 12.5 10.1 8.8 8.2 20.3
+According to Table 6, multiple iterations can significantly reduce quantization error.
+For instance, using QPiSSA-r64 with 5-iter on LLaMA-3-8B reduces the quantization error nearly twice as much as with 1-iter.
+In the main paper, we used 5 iterations in Section 5.3 and Section 5.4, while 1 iteration was used in Section 5.5.
+Appendix F Comparing the Quantization Error of QLoRA, LoftQ and QPiSSA
+This section extends the discussion in Section 4 by providing a comprehensive comparison of the quantization errors associated with QLoRA, LoftQ, and QPiSSA.
+Using the “layers[0].self_attn.q_proj” of LLaMA 2-7B as an example, we illustrate the singular values of critical matrices during the quantization process with QLoRA, LoftQ, and PiSSA in Figure 9.
+A larger sum of the singular values (nuclear norm) of the error matrix indicates a greater quantization error.
+(a) Original matrix $W$
+(b) Quantized matrix $nf4(W)$
+(c) Residual matrix $W^{res}$
+(d) Error of QLoRA
+(e) Error of LoftQ
+(f) Error of PiSSA
+Figure 9: Several important singular values for calculation the quantization error of QLoRA, LoftQ and PiSSA.
+The quantization error of QLoRA, which quantizes the base model to Normal Float 4-bit (NF4) and initializes $A$ and $B$ with Gaussian-Zero initialization, is:
+Quantization Error of $\text{Quantization Error of QLoRA}=||W-\left(nf4(W)+AB\right)||_{*}=||W-nf4(W)||_{*},$ (13)
+As shown in Equation 13, QLoRA decomposes the original matrix in Figure 8(a) into the sum of a quantized matrix (Figure 8(b)) and an error matrix (Figure 8(d)).
+By comparing Figure 8(a) and Figure 8(d), we can see that the magnitude of the error matrix is much smaller than that of the original matrix.
+Therefore, the benefit of preserving the principal components of the $W$ matrix with the adapter is greater than that of preserving the principal components of the error matrix with the adapter.
+LoftQ [49], designed to preserve the principal components of the error matrix using the adapter, first performs singular value decomposition on the quantization error matrix of QLoRA:
+$U^{err}S^{err}V^{err}=W-nf4(W),$ (14)
+then uses the larger singular values to initialize $A$ and $B$, thereby reducing the quantization error to:
+$LoftQ^{err}=||W-\left(nf4(W)+AB\right)||*=||U^{err}_{[r:]}S^{err}_{[r:,r:]}V^{err}_{[r:]}||*=\sum_{i=r}^{min(m,n)}{S^{err}_{[i,i]}}.$ (15)
+LoftQ eliminates only the largest $r$ singular values $S^{\text{err}}_{[:r]}$ (see Figure 8(e)) from the QLoRA error matrix (Figure 8(d)).
+Our PiSSA, however, does not quantify the base model but the residual model:
+Quantization Error of $\text{Quantization Error of PiSSA}=||W-\left(nf4(W^{res})+AB\right)||_{*}=||W^{res}-nf4(W^{res})||_{*},$ (16)
+where $A$ and $B$ are initialized following Equation 2 and 3.
+Since the residual model has removed the large-singular-value components, the value distribution of $W^{res}$ can be better fitted by a Student’s t-distribution with higher degrees of freedom compared to $W$ (as can be seen in Figure 10) and thus quantizing $W^{res}$ results in lower error using 4-bit NormalFloat (shown in Figure 8(f)).
+(a) The original matrix $W$
+(b) The residual matrix $W_{r}es$
+Figure 10: Fitting the original matrix and the residual matrix using Student’s t-distribution.
+Appendix G Evaluating PiSSA on Mixtral and Gemma with More Training Steps
+This is the supplement for Section 5.2.
+We applied PiSSA, LoRA, and full parameter fine-tuning on the full MetaMathQA-395K dataset using Mistral-7B and Gemma-7B models, training for 3 epochs.
+Figures 11 and 12 display the training loss, gradient norm, and evaluation accuracy on GSM8K.
+(a) Loss over training steps.
+(b) Grad norm over steps.
+(c) Accuracy on GSM8K over training steps.
+Figure 11: Fine-tuning Mistral-7B-v0.1 on the MetaMathQA-395K dataset for 3 epochs: A comparison of full parameter fine-tuning (indicated by a dashed line), LoRA (in blue), and PiSSA (in orange).
+(a) Loss over training steps.
+(b) Grad norm over steps.
+(c) Accuracy on GSM8K over training steps.
+Figure 12: Fine-tuning Gemma-7B on the MetaMathQA-395K dataset for 3 epochs: A comparison of full parameter fine-tuning (indicated by a dashed line), LoRA (in blue), and PiSSA (in orange).
+As shown in Figure 10(a) and 11(a), the loss for full parameter fine-tuning decreases sharply with each epoch, indicating overfitting to the training data.
+Notably, during the entire first epoch, the loss for full parameter fine-tuning on Mistral and Gemma is significantly higher than for LoRA and PiSSA, suggesting that full parameter fine-tuning has weaker generalization capabilities compared to LoRA and PiSSA on Mistral-7B and Gemma-7B models.
+The gradient norm for the first epoch in Figure 11(b) fluctuates dramatically with each step, further indicating instability in the training process for full parameter fine-tuning.
+Consequently, as illustrated in Figures 10(c) and 11(c), the performance of full parameter fine-tuning is markedly inferior to that of LoRA and PiSSA.
+These experiments demonstrate that using parameter-efficient fine-tuning can prevent the over-fitting issue caused by over-parameters.
+Appendix H Supplementary Experiments on Various Ranks
+H.1 Quantization Error for More Type of Layers
+Figure 7(a) only shows the reduction ratio of quantization error for “q_proj” layers.
+In Figure 13, we present the error reduction ratios for the remaining types of linear layers under different ranks.
 Figure 13: Comparison of quantization errors in QLoRA, LoftQ, and PiSSA across k_proj, v_proj, o_proj and gate_proj, up_proj, down_proj layers.
+From Figure 13 it can be observed that under different ranks, the reduction ratio of quantization error for various linear layers in LLaMA-2-7B, including “k_proj”, “v_proj”, “o_proj”, “gate_proj”, “up_proj”, and “down_proj” layers, is consistently lower with PiSSA compared to LotfQ.
+H.2 Evaluation Performance for More Model on Various Ranks
+Section 5.5 only validated the effectiveness of LLaMA-2-7B. In Figure 14, we also present the comparative results of Mistral-7B-v0.1, and Gemma-7B under different ranks.
+(a) Final training loss across different ranks.
+(b) Rank-wise performance evaluated using pass@1 on the GSM8K dataset.
+(c) Rank-wise performance evaluated using pass@1 on the MATH dataset.
+Figure 14: Fine-tuning LLaMA 2-7B, Mistral-7B-v0.1, and Gemma-7B on the MetaMathQA dataset: A comparison of full parameter fine-tuning (indicated by a dashed line), LoRA (in blue), and PiSSA (in orange).
+From Figure 14, PiSSA uses fewer trainable parameters compared to LoRA while achieving or even surpassing full-parameter fine-tuning on LLaMA-2-7B and Mistral-7B.
+Remarkably, on Gemma-7B, PiSSA exceeds full-parameter fine-tuning performance even at rank=1.
+However, as the rank increases to 128, the performance of PiSSA begins to decline, indicating that PiSSA over-parameterizes earlier than LoRA.
+This over-parameterization phenomenon does not occur on LLaMA-2-7B, suggesting that increasing the rank further might enable PiSSA to achieve even higher performance on LLaMA-2-7B.
+H.3 More Training Loss and Grad Norm under Various Ranks
+In Figure 15 and 16, we examining the loss and gradient norm during the training process of PiSSA and LoRA on LLaMA 2-7B, Mistral-7B-v0.1, and Gemma-7B using different ranks.
+Figure 15: Comparison of training loss for LLaMA-2-7B, Mistral-7B, and Gemma-7B, organized into three rows, using LoRA and PISSA across ranks $2^{i},i\in[0,7]$, organized into eight columns.
+Figure 16: Comparison of grad norm for LLaMA-2-7B, Mistral-7B, and Gemma-7B, organized into three rows, using LoRA and PISSA across ranks $2^{i},i\in[0,7]$, organized into eight columns.
+From Figure 15, PiSSA consistently shows a faster initial loss reduction compared to LoRA across various ranks.
+Additionally, the final loss remains lower than that of LoRA.
+This advantage is particularly pronounced when the rank is smaller.
+From Figure 16, the gradient norm of PiSSA remains consistently higher than that of LoRA throughout the training process, indicating its efficient fitting of the training data.
+A closer look at the first few steps of LoRA’s gradient norm reveals a trend of rising and then falling.
+According to Section 3, LoRA’s gradients are initially close to zero, leading to very slow model updates.
+This requires several steps to elevate LoRA’s weights to a higher level before subsequent updates.
+This phenomenon validates our assertion that LoRA wastes some training steps and therefore converges more slowly.
+It demonstrates the robustness of the faster convergence property of PiSSA across various ranks.
+Appendix I Experimental Settings on NLU
+Datasets
+We evaluate the performance of PiSSA on GLUE benchmark, including 2 single-sentence classification tasks (CoLA, SST), 5 pairwise text classification tasks (MNLI, RTE, QQP, MRPC and QNLI) and 1 text similarity prediction task (STS-B).
+We report overall matched and mismatched accuracy on MNLI, Matthew’s correlation on CoLA, Pearson correlation on STS-B, and accuracy on the other datasets.
+Implementation Details
+To evaluate the performance of PiSSA intuitively, we compared PiSSA and LoRA with the same number of trainable parameters.
+RoBERTa-large has $355M$ trainable parameters.
+PiSSA was applied to $W_{Q}$ and $W_{V}$, resulting in a total of $1.84M$ trainable parameters.
+DeBERTa-v3-base has $184M$ trainable parameters.
+PiSSA and LoRA were applied to $W_{Q}$, $W_{K}$ and $W_{V}$ respectively, resulting in a total of $1.33M$ trainable parameters.
+Experiments on NLU are based on the publicly available LoftQ [49] code-base.
+On RoBERTa-large, we initialize the model with pretrained MNLI checkpoint for the task of MRPC, RTE and STS-B.
+We set the rank of PiSSA in this experiment as $8$ and selecte lora alpha in 8, 16.
+We utilize AdamW with linear learning rate schedule to optimize and tune learning rate (LR) from 1e-4,2e-4,3e-4,4e-4,5e-4, 6e-4, 5e-5, 3e-5.
+Batch sizes (BS) are selected from ${6,8,16,32}$.
+Table 7 presents the hyperparmeters we utilized on the glue benchmark.
+Table 7: Hyperparameters of our Methods on GLUE.
+Dataset Roberta-large Deberta-v3-base
+Epoch BS LR Lora alpha Epoch BS LR Lora alpha
+MNLI 10 32 1e-4 16 5 16 5e-5 8
+SST-2 10 32 2e-4 16 20 16 3e-5 8
+MRPC 20 16 6e-4 8 20 32 2e-4 8
+CoLA 20 16 4e-4 8 20 16 1e-4 8
+QNLI 10 6 1e-4 8 10 32 1e-4 16
+QQP 20 32 3e-4 8 10 16 1e-4 8
+RTE 20 16 3e-4 16 50 16 1e-4 8
+STS-B 30 16 3e-4 16 20 8 3e-4 8

### v4

@@
-Fanxu Meng1,2,3, Zhaohui Wang1, Muhan Zhang1,3
+Fanxu Meng1,2, Zhaohui Wang1, Muhan Zhang1,2
 Abstract
@@
 Compared to LoRA, PiSSA updates the principal components while freezing the “residual” parts, allowing faster convergence and enhanced performance.
-Comparative experiments of PiSSA and LoRA across 12 different models, ranging from 184M to 70B, encompassing 5 NLG and 8 NLU tasks, reveal that PiSSA consistently outperforms LoRA under identical experimental setups.
-On the GSM8K benchmark, Mistral-7B fine-tuned with PiSSA achieves an accuracy of 72.86%, surpassing LoRA’s 67.7% by 5.16%.
+Comparative experiments of PiSSA and LoRA across 11 different models, ranging from 184M to 70B, encompassing 5 NLG and 8 NLU tasks, reveal that PiSSA consistently outperforms LoRA under identical experimental setups.
+On the GSM8K benchmark, Gemma-7B fine-tuned with PiSSA achieves an accuracy of 77.7%, surpassing LoRA’s 74.53% by 3.25%.
 Due to the same architecture, PiSSA is also compatible with quantization to further reduce the memory requirement of fine-tuning.
 Compared to QLoRA, QPiSSA (PiSSA with 4-bit quantization) exhibits smaller quantization errors in the initial stages.
-Fine-tuning LLaMA-3-70B on GSM8K, QPiSSA attains an accuracy of 86.05%, exceeding the performances of QLoRA at 81.73%.
+Fine-tuning LLaMA-3-70B on GSM8K, QPiSSA attains an accuracy of 86.05%, exceeding the performance of QLoRA at 81.73%.
 Leveraging a fast SVD technique, PiSSA can be initialized in only a few seconds, presenting a negligible cost for transitioning from LoRA to PiSSA.
@@
 Figure 1: The comparison among Full Fine-tuning, training with LoRA, and PiSSA.
-In this visualization, blue modules represent parts of the model where parameters are frozen during training, while orange modules indicate components that require updates.
+In this visualization, blue modules represent parts of the model whose parameters are frozen during training, while orange modules indicate components that require updates.
 QLoRA quantizes the pretrained matrix in LoRA to 4-bit, whereas QPiSSA quantizes the residual matrix in PiSSA.
-(a) PiSSA converges more rapidly.
-(b) PiSSA reduces more quantization error.
-Figure 2: We illustrate the two key advantages of PiSSA: converging faster and better, and reducing quantization error.
-In the left figure, we use a toy example to show PiSSA’s faster convergence, where we first train a two-layer MLP classifying odd numbers of MNIST, and then fine-tune the model on even numbers.
-PiSSA finds the right direction more quickly and achieves a lower loss with the same number of steps.
-In the right figure, we show the quantization error reduction of PiSSA compared to directly quantizing the base model.
-PiSSA can reduce more quantization error than the baselines.
+Table 1: Comparison of similarities and differences between PiSSA and LoRA. In this table, bold highlights the model’s primary component, while underline denotes the residual component.
+LoRA PiSSA
+Forward $Y=X(\textbf{W}+\underline{\Delta W})=X(\textbf{W}+\underline{AB})$ $Y=X(\underline{W^{res}}+\textbf{W}^{\textbf{pri}})=X(\underline{W^{\text{res}}}+\textbf{AB})$
+$\underline{A}\sim\mathcal{N}(0,\sigma^{2})\in\mathbb{R}^{m\times r}$ $\textbf{A}=U_{[:,\textbf{:r}]}\,S_{[\textbf{:r},\textbf{:r}]}^{1/2}\in\mathbb{R}^{m\times r}$
+Initialization $\underline{B}=0\in\mathbb{R}^{r\times n}$ $\textbf{B}=S_{[\textbf{:r},\textbf{:r}]}^{1/2}\,V_{[:,\textbf{:r}]}^{\top}\in\mathbb{R}^{r\times n}$
+$\underline{W^{\text{res}}}=U_{[:,\underline{r:}]}\,S_{[\underline{r:},\underline{r:}]}\,V_{[:,\underline{r:}]}^{\top}\in\mathbb{R}^{m\times n}$
+Gradient $\frac{\partial L}{\partial A}=X^{\top}\left(\frac{\partial L}{\partial Y}\right)\underline{B}^{\top}\rightarrow\underline{0}$ $\frac{\partial L}{\partial A}=X^{\top}\left(\frac{\partial L}{\partial Y}\right)\textbf{B}^{\top}\rightarrow\textbf{Principal}$
+$\frac{\partial L}{\partial B}=\underline{A^{\top}}X^{\top}\left(\frac{\partial L}{\partial Y}\right)\rightarrow$ Random Direction $\frac{\partial L}{\partial B}=\textbf{A}^{\top}X^{\top}\left(\frac{\partial L}{\partial Y}\right)\rightarrow\textbf{Principal}$
+Comparison Fine-tunes noise while freezing W. Fine-tunes principal parts freezing $W^{\text{res}}$.
+Slow convergence and underperformance. Fast convergence and better performance.
+QLoRA cannot reduce quantization error. QPiSSA can reduce quantization error.
 1 Introduction
-Fine-tuning large language models (LLMs) is a highly effective technique for boosting their capabilities in various tasks [1, 2, 3, 4], ensuring models to follow instructions [5, 6, 7], and instilling models with desirable behaviors while eliminating undesirable ones [8, 9].
+Fine-tuning large language models (LLMs) is a highly effective technique for boosting their capabilities in various tasks luo2023wizardmath ; yu2023metamath ; luo2023wizardcoder ; li2023starcoder , ensuring models to follow instructions ouyang2022training ; zheng2024judging ; xu2023wizardlm , and instilling models with desirable behaviors while eliminating undesirable ones bai2022training ; rafailov2024direct .
 However, the fine-tuning process for very large models is accompanied by prohibitive costs.
-For example, regular 16-bit fine-tuning of a LLaMA 65B parameter model requires over 780 GB of GPU memory [10], and the VRAM consumption for training GPT-3 175B reaches 1.2TB [11].
-Consequently, various parameter-efficient fine-tuning (PEFT) [12, 13] methods have been proposed to reduce the number of parameters and memory usage required for fine-tuning.
-Due to the ability to maintain the performance of full fine-tuning without adding additional inference latency, Low-Rank Adaptation (LoRA) [11] has emerged as a popular PEFT method.
-LoRA [11] hypothesizes that the modifications to parameter matrices during fine-tuning exhibit low-rank properties.
+For example, regular 16-bit fine-tuning of a LLaMA 65B parameter model requires over 780 GB of GPU memory dettmers2024qlora , and the VRAM consumption for training GPT-3 175B reaches 1.2TB hu2021lora .
+Consequently, various parameter-efficient fine-tuning (PEFT) xu2023parameter ; han2024parameter methods have been proposed to reduce the number of parameters and memory usage required for fine-tuning.
+Due to the ability to maintain the performance of full fine-tuning without adding additional inference latency, Low-Rank Adaptation (LoRA) hu2021lora has emerged as a popular PEFT method.
+LoRA hu2021lora hypothesizes that the modifications to parameter matrices during fine-tuning exhibit low-rank properties.
 As depicted in Figure 1(b), for a pre-trained weight matrix $W\in\mathbb{R}^{m\times n}$, LoRA substitutes the updates with a low-rank decomposition $\Delta W=AB$, where $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$, and the rank $r\ll\text{min}(m,n)$.
@@
 $Y=X(W+\Delta W)=X(W+AB),$ (1)
-where $X\in\mathbb{R}^{b\times m}$, $Y\in\mathbb{R}^{b\times n}$, and $b$ represents the batch size of input data.
 A random Gaussian initialization is used for $A$ and zero for $B$, making $AB=0$ at the beginning of training, thereby the injection of adapters does not affect the model’s output initially.
 LoRA avoids the need to compute gradients or maintain the optimizer states for the original matrix $W$, instead optimizing the injected, significantly smaller low-rank matrices $A,B$.
-Thus, it could reduce the number of trainable parameters by 10,000 times and the GPU memory requirement by 3 times [11].
-Moreover, LoRA often achieves comparable or superior performance to full parameter fine-tuning, indicating that fine-tuning “parts” of the full parameters can be enough for downstream tasks.
-By integrating the quantization of pre-trained matrices $W$, LoRA also enables reducing the average memory requirements by 16 times [10].
-Meanwhile, the adapters are still allowed to utilize higher precision weights; thus, the quantization usually does not significantly degrade the performance of LoRA.
-According to Equation 1, the gradients of A and B are $\frac{\partial L}{\partial A}=X^{T}\left(\frac{\partial L}{\partial Y}\right)B^{T}$ and $\frac{\partial L}{\partial B}=A^{T}X^{T}\left(\frac{\partial L}{\partial Y}\right)$.
-Compared to full fine-tuning, using LoRA initially does not change the output $Y$ for the same input $X$, so the gradient magnitude is primarily determined by the values of $A$ and $B$.
-Since $A$ and $B$ are initialized with Gaussian noise and zeros in LoRA, the gradients can be very small, leading to slow convergence in the fine-tuning process.
-We also observe this phenomenon empirically, where LoRA often wastes much time around the initial point.
+Thus, it could reduce the number of trainable parameters by 10,000$\times$ and the GPU memory requirement by 3$\times$ hu2021lora .
+LoRA is capable of achieving comparable performance to full parameter fine-tuning.
+By integrating the quantization of pre-trained matrices $W$, LoRA also enables reducing the average memory requirements by 16$\times$ dettmers2024qlora .
+Meanwhile, the adapters can still utilize higher precision weights, thus, the quantization usually does not significantly degrade the performance of LoRA.
+According to Equation 1, the gradients of A and B are $\frac{\partial L}{\partial A}=X^{\top}\left(\frac{\partial L}{\partial Y}\right)B^{\top}$ and $\frac{\partial L}{\partial B}=A^{\top}X^{\top}\left(\frac{\partial L}{\partial Y}\right)$.
+Compared to full fine-tuning, using LoRA initially does not change the output $Y$ for the same input $X$, so the magnitude and direction of gradient are primarily determined by the values of $A$ and $B$.
+Since $A$ and $B$ are initialized with Gaussian noise and zeros in LoRA, the gradients could be small and uninformative for a long time, leading to slow convergence in the fine-tuning process.
+We also observe this phenomenon empirically, as LoRA often wastes much time around the initial point.
 Our Principal Singular values and Singular vectors Adapter (PiSSA) diverges from LoRA and its successors by focusing not on approximating $\Delta W$, but $W$.
@@
 Moreover, the loss and gradient norm curves of PiSSA often demonstrate a similar trend to those of full parameter fine-tuning in our experiments (Figure 4), indicating that fine-tuning the principal components matches the behavior of fine-tuning the full matrix to some degree.
-Since the principal components $W^{pri}$ are preserved in the adapter and are of full precision, another benefit of PiSSA is that when applying quantization to the frozen part $W^{res}$, we can significantly reduce the quantization error compared to QLoRA (which quantizes the whole $W$), as illustrated in Figure 2(b).
-Therefore, PiSSA is perfectly compatible with quantization, making it a plug-and-play substitution for LoRA.
+(a) PiSSA converges more rapidly.
+(b) PiSSA reduces more quantization error.
+Figure 2: We illustrate the two key advantages of PiSSA: converging faster and better, and reducing quantization error.
+In the left figure, we use a toy example to show PiSSA’s faster convergence, where we first train a two-layer MLP classifying odd numbers of MNIST, and then fine-tune the model on even numbers.
+PiSSA finds the right direction more quickly and achieves a lower loss with the same number of steps.
+In the right figure, PiSSA reduces quantization error more effectively than LoftQ li2023loftq , with an optional 5-iteration SVD for further error reduction, as detailed in Appendix E.
+Because the principal components $W^{pri}$ are preserved in the adapter at full precision, an additional benefit of PiSSA is that when applying quantization to the frozen part $W^{res}$, we can significantly reduce the quantization error compared to QLoRA (which quantizes the entire $W$), as illustrated in Figure 2(b).
+Therefore, PiSSA is even more compatible with quantization than LoRA, making it a superior plug-and-play substitution for LoRA.
+Our paper makes several significant contributions:
+  • We analyze the initial gradient magnitude and direction in LoRA, demonstrating that $A$ initially has a zero gradient and $B$ has a random gradient, which slows down convergence and may lead to convergence at suboptimal local minima.
+  • We propose PiSSA initialization, a novel method that approximates the optimization direction of full-parameter fine-tuning by adapting a model’s principal components.
+To our knowledge, PiSSA is the first to apply SVD to the original model, using principal singular values and vectors to initialize the adapter for fine-tuning, while keeping the residual components frozen.
+Experiments show that PiSSA converges faster and outperforms LoRA.
+  • We combine PiSSA with NF4 quantization to propose QPiSSA, which reduces quantization error by about 20% compared to QLoRA, while maintaining the fast convergence and high performance of PiSSA.
 2 Related Works
 The vast complexity and computational needs of large language models (LLMs) with billions of parameters present significant hurdles in adapting them for specific downstream tasks.
-Parameter Efficient Fine-Tuning (PEFT) [12, 13] emerges as a compelling solution by minimizing the fine-tuning parameters and memory requirements while achieving comparable performance to full fine-tuning.
-PEFT encompasses strategies like partial fine-tuning [14, 15, 16, 17, 18, 19, 20, 21], soft prompt fine-tuning  [22, 23, 24, 25, 26, 27, 28], non-linear adapter fine-tuning [29, 30, 31, 32, 33, 34, 35, 36, 37, 38], and low rank adapter based fine-tuning [39, 40, 11, 41].
-LoRA [11] inject trainable adapters to the linear layers.
+Parameter Efficient Fine-Tuning (PEFT) xu2023parameter ; han2024parameter emerges as a compelling solution by minimizing the fine-tuning parameters and memory requirements while achieving comparable performance to full fine-tuning.
+PEFT encompasses strategies like partial fine-tuning zaken2021bitfit ; lawton2023neural ; zhao2020masking ; sung2021training ; ansell2021composable ; xu2021raise ; guo2020parameter ; fu2023effectiveness , soft prompt fine-tuning  hambardzumyan2021warp ; lester2021power ; li2021prefix ; liu2023gpt ; vu2021spot ; asai2022attempt ; wang2023multitask , non-linear adapter fine-tuning houlsby2019parameter ; lin2020exploring ; lei2024conditional ; he2021towards ; ruckle2020adapterdrop ; zhao2022tiny ; pfeiffer2020adapterfusion ; he2023mera ; mahabadi2021parameter ; chronopoulou2023adaptersoup , and low rank adapter based fine-tuning li2018measuring ; aghajanyan2020intrinsic ; hu2021lora ; zhang2022adaptive .
+LoRA hu2021lora injects trainable adapters to the linear layers.
 After fine-tuning, these adaptations can be re-parameterized into the standard model structure, thus gaining widespread adoption due to their ability to maintain the model’s original architecture while enabling efficient fine-tuning.
-Following LoRA, AdaLoRA [42, 41, 43] dynamically learns the rank size needed for LoRA in each layer of the model.
-DeltaLoRA [44, 45] updates the original weights of the model using parameters from adapter layers, enhancing LoRA’s representational capacity.
-LoSparse [46] incorporates LoRA to prevent pruning from eliminating too many expressive neurons.
-DoRA [47] introduces a magnitude component to learn the scale of $\Delta W$ while utilizing the original AB as a direction component of $\Delta W$.
-Unlike LoRA and its successors, which focus on learning low-rank approximations of weight updates, our PiSSA approach directly tunes the essential but low-rank parts of the model while keeping the noisier, high-rank, and nonessential parts frozen.
+Following LoRA, AdaLoRA zhang2022adaptive dynamically learns the rank size needed for LoRA in each layer of the model.
+DeltaLoRA zi2023delta updates the original weights of the model using parameters from adapter layers, enhancing LoRA’s representational capacity.
+LoSparse li2023losparse incorporates LoRA to prevent pruning from eliminating too many expressive neurons.
+DoRA liu2024dora introduces a magnitude component to learn the scale of $\Delta W$ while utilizing the original AB as a direction component of $\Delta W$.
+Unlike LoRA and its successors, which focus on learning low-rank approximations of weight updates, our PiSSA directly tunes the essential low-rank parts of the model while keeping the noisier, high-rank, and nonessential parts frozen.
 Although our approach differs in philosophy from LoRA, it shares most of LoRA’s structural benefits and can be extended by these methods to enhance its performance.
-QLoRA [10] integrates LoRA with 4-bit NormalFloat (NF4) quantization, along with Double Quantization and Paged Optimizers, enabling the fine-tuning of a 65B parameter model on a single 48GB GPU while preserving the performance of full 16-bit fine-tuning tasks.
-QA-LoRA [48] introduces group-wise operators to increase the degree of freedom in low-bit quantization.
-LoftQ [49] reduces quantization error by decomposing the quantization error matrix of QLoRA and retaining the principal components with an adapter.
-Our PiSSA approach can also be combined with quantization techniques, and we have found that PiSSA significantly reduces quantization error compared to QLoRA and LoftQ.
+QLoRA dettmers2024qlora integrates LoRA with 4-bit NormalFloat (NF4) quantization, along with Double Quantization and Paged Optimizers, enabling the fine-tuning of a 65B parameter model on a single 48GB GPU while preserving the performance of full 16-bit fine-tuning tasks.
+QA-LoRA xu2023qa introduces group-wise operators to increase the degree of freedom in low-bit quantization.
+LoftQ li2023loftq reduces quantization error by decomposing the quantization error matrix of QLoRA and retaining the principal components with an adapter.
+PiSSA can also be combined with quantization techniques, and we have found that PiSSA significantly reduces quantization error compared to QLoRA and LoftQ.
 3 PiSSA: Principal Singular Values and Singular Vectors Adaptation
@@
 $A$ $=U_{[:,:r]}\,S_{[:r,:r]}^{1/2}\in\mathbb{R}^{m\times r},$ (2)
-$B$ $=S_{[:r,:r]}^{1/2}\,V_{[:,:r]}^{T}\in\mathbb{R}^{r\times n}.$ (3)
+$B$ $=S_{[:r,:r]}^{1/2}\,V_{[:,:r]}^{\top}\in\mathbb{R}^{r\times n}.$ (3)
 The residual singular values and vectors are used to build the residual matrix which is frozen during fine-tuning:
-$W^{res}=U_{[:,r:]}\,S_{[r:,r:]}\,V_{[:,r:]}^{T}\in\mathbb{R}^{m\times n}.$ (4)
+$W^{res}=U_{[:,r:]}\,S_{[r:,r:]}\,V_{[:,r:]}^{\top}\in\mathbb{R}^{m\times n}.$ (4)
 As indicated by Equation 5, the integration of $AB$ with the residual matrix also preserves the full capability of the pre-trained model in the beginning of fine-tuning:
 $Y=XW=X(W^{res}+W^{pri})=X(W^{res}+AB).$ (5)
-Similar to LoRA, the gradients of $A$ and $B$ are also given by $\frac{\partial L}{\partial A}=X^{T}\left(\frac{\partial L}{\partial Y}\right)B^{T}$ and $\frac{\partial L}{\partial B}=A^{T}X^{T}\left(\frac{\partial L}{\partial Y}\right)$.
+Similar to LoRA, the gradients of $A$ and $B$ are also given by $\frac{\partial L}{\partial A}=X^{\top}\left(\frac{\partial L}{\partial Y}\right)B^{\top}$ and $\frac{\partial L}{\partial B}=A^{\top}X^{\top}\left(\frac{\partial L}{\partial Y}\right)$.
 Since elements of $\mathbf{s}_{[:r]}$ $\gg$ elements of $\mathbf{s}_{[r:]}$, the trainable adapter $W^{pri}=AB$ contains the most essential directions of $W$.
@@
 Consequently, the gradients are small or in random directions during the early stages of fine-tuning, possibly introducing much waste of gradient descent steps.
-Moreover, an inferior initialization might lead to suboptimal local minimum points found, causing worse generalization performance.
+Moreover, an inferior initialization might lead to suboptimal local minimum, resulting in worse generalization performance.
 Since PiSSA shares the identical architecture with LoRA, it inherits most of LoRA’s benefits.
@@
 The adapter’s straightforward linear structure facilitates the integration of trainable matrices with the pre-trained weights upon deployment, thereby maintaining the original inference speed of a fully fine-tuned model.
-Employing the Fast SVD technique [50] allowed PiSSA to finish initialization in several seconds (Appendix B), which is a negligible cost.
+Employing the Fast SVD technique halko2011finding allowed PiSSA to finish initialization in several seconds (Appendix B), which is a negligible cost.
 For storage efficiency, we can choose not to store the dense parameter matrix $\Delta W$, but to store the low-rank matrices, $\Delta A$ and $\Delta B$ instead.
@@
 Finally, one pre-trained model can accommodate multiple $\Delta A,\Delta B$, fine-tuned by diverse PiSSA or LoRA procedures, which enables fast adaptation of the pre-trained model to different downstream applications.
-4 QPiSSA: PiSSA with Quantization
+4 QPiSSA: An Extension Method with Lower Quantization Error
 Quantization divides the value range of a matrix into several continuous regions, and maps all values falling inside a region into the same “quantized” value.
-It is an effective technique to reduce the memory consumption of forward propagation, but breaks down during backpropagation.
+It is an effective technique to reduce the memory consumption of forward propagation.
 At the same time, LoRA greatly reduces the backward memory requirement, making it highly suitable to use LoRA and quantization together, where the base model is quantized for memory-efficient forward propagation, and the LoRA adaptors are kept in full precision for accurate backward parameter updates.
@@
 Quantization Error of $\text{Quantization Error of QLoRA}=||W-\left(nf4(W)+AB\right)||_{*}=||W-nf4(W)||_{*},$ (6)
-where $||M||_{*}$ denotes the nuclear norm (also known as the trace norm)  [51], defined as:
+where $||M||_{*}$ denotes the nuclear norm (also known as the trace norm fan1951maximum ), defined as:
 $\|M\|_{*}=\operatorname{trace}\left({\sqrt{M^{*}M}}\right)=\sum_{i=1}^{\min\{m,n\}}\sigma_{i}(M),$ (7)
@@
 Since the residual model has removed the large-singular-value components, $W^{res}$ has a narrower distribution than that of $W$, as can be seen in Figures 3(a) and 3(b) (comparing the singular value distributions of $W$ and $W^{res}$), as well as Figures 3(c) and 3(f) (comparing the value distributions of $W$ and $W^{res}$), which is beneficial for reducing the quantization error.
-Furthermore, since NF4 is optimized for normally distributed data [10], we fit a Gaussian for the values of $W$ and $W^{res}$, respectively.
-As can be seen in Figures 3(c) and 3(f), $W^{res}$ is much more Gaussian-like and has a smaller standard deviation, making it more suitable to apply NF4 to $W^{res}$ instead of $W$.
+Moreover, given that NF4 is optimized for data with a normal distribution, as discussed by Dettmers et al.
+dettmers2024qlora , we fit the values of $W$ and $W^{res}$ to a Gaussian distribution respectively.
+As illustrated in Figures 3(c) and 3(f), $W^{res}$ aligns more closely with a Gaussian distribution and exhibits a smaller standard deviation, making it more suitable for applying NF4 than $W$.
 Both the above lead QPiSSA to achieve a significantly lower quantization error than QLoRA, shown in Figures 3(d) and 3(e).
@@
 (f) The distribution of $W^{res}$
-Figure 3: Figures (a), (b), (d) and (e) visualize the singular values in descending order for $W$, $W^{res}$, $W-nf4(W)$ and $W^{res}-nf4(W^{res})$ of LLaMA 2-7B’s “layers[0].self_attn.q_proj” matrix.
-Figures (c) and (f) show the data value distributions of $W$ and $W^{res}$.
+Figure 3: Visualizations of LLaMA 2-7B’s “layers[0].self_attn.q_proj” matrix, with distributions for the full model shown in Appendix G.
+Figures (a), (b), (d), and (e) display the singular values of $W$, $W^{res}$, $W-nf4(W)$, and $W^{res}-nf4(W^{res})$, respectively.
+Figures (c) and (f) show the data distributions of $W$ and $W^{res}$.
 Besides the advantage of reducing quantization error, QPiSSA’s gradient direction is similar to that of PiSSA, resulting in significantly better fine-tuning performance compared to QLoRA.
@@
 The experiments were conducted on the NVIDIA A800-SXM4(80G) GPU.
-In our experiments, we adopt the Alpaca [52] implementation strategy, using the AdamW optimizer with a batch size of 128, a learning rate of 2e-5, cosine annealing schedules, and a warmup ratio of 0.03, without any weight decay.
-As discussed in Section B.3 of QLoRA [10], we compute the loss using only the responses from the instruction-following datasets.
+In our experiments, we adopt the Alpaca alpaca implementation strategy, using the AdamW optimizer with a batch size of 128, a learning rate of 2e-5, cosine annealing schedules, and a warmup ratio of 0.03, without any weight decay.
+As discussed in Section B.3 of QLoRA dettmers2024qlora , we compute the loss using only the responses from the instruction-following datasets.
 We ensure lora_alpha is always equal to lora_r, set lora_dropout to 0, and incorporate the adapters into all linear layers of the base model.
 We utilize the Float32 computation type for both the base model and the adapter in LoRA and PiSSA.
-For QLoRA, LoftQ, and QPiSSA, we use 4-bit NormalFloat [10] for the base model and Float32 for the adapter.
-BFloat16 [53] is used for full parameter fine-tuning to save the resources (see Appendix D).
+For QLoRA, LoftQ, and QPiSSA, we use 4-bit NormalFloat dettmers2024qlora for the base model and Float32 for the adapter.
+BFloat16 wang2019bfloat16 is used for full parameter fine-tuning to save the resources (see Appendix D).
 5.1 Evaluating the Performance of PiSSA on both NLG and NLU Tasks
-Table 1 presents a comparative evaluation of fine-tuning strategies PiSSA, LoRA, and full parameter fine-tuning on natural language generation (NLG) tasks.
-We fine-tuned LLaMA 2-7B [54], Mistral-7B-v0.1 [55], and Gemma-7B [56] on the MetaMathQA dataset [2] to assess their mathematical problem-solving capabilities on the GSM8K [57] and MATH [2] validation sets.
-Additionally, the models were fine-tuned on the CodeFeedback dataset [58] and evaluated for coding proficiency using the HumanEval [59] and MBPP [60] datasets.
-Furthermore, the models were trained on the WizardLM-Evol-Instruct dataset [7] and tested for conversational abilities on the MT-Bench dataset [6].
+We begin by comparing PiSSA, LoRA, and full-parameter fine-tuning on natural language generation (NLG) tasks.
+We fine-tuned LLaMA 2-7B touvron2023llama , Mistral-7B-v0.1 jiang2023mistral , and Gemma-7B team2024gemma on the MetaMathQA dataset yu2023metamath to assess their mathematical problem-solving capabilities on the GSM8K cobbe2021gsm8k and MATH hendrycks2021measuring validation sets.
+Additionally, the models were fine-tuned on the CodeFeedback dataset zheng2024opencodeinterpreter and evaluated for coding proficiency using the HumanEval chen2021evaluating and MBPP austin2021program datasets.
+Furthermore, the models were trained on the WizardLM-Evol-Instruct dataset xu2023wizardlm and tested for conversational abilities on the MT-Bench dataset zheng2024judging .
 All experiments were conducted using subsets containing 100K data points and were trained for only one epoch to reduce training overhead.
-Table 1: Comparing PiSSA and LoRA on NLG tasks
-Model Strategy Trainable GSM8K MATH HumanEval MBPP MT-Bench
-Parameters
-LLaMA 2-7B Full FT 6738M 49.05 7.22 21.34 35.59 4.91
-LoRA 320M 42.30 5.50 18.29 35.34 4.58
-PiSSA 320M 53.07 7.44 21.95 37.09 4.87
-Mistral-7B Full FT 7242M 67.02 18.6 45.12 51.38 4.95
-LoRA 168M 67.70 19.68 43.90 58.39 4.90
-PiSSA 168M 72.86 21.54 46.95 62.66 5.34
-Gemma-7B Full FT 8538M 71.34 22.74 46.95 55.64 5.40
-LoRA 200M 74.90 31.28 53.66 65.41 4.98
-PiSSA 200M 77.94 31.94 54.27 66.17 5.64
+Table 2: Comparison of PiSSA and LoRA on NLG tasks, with results averaged over three runs and reported with standard deviations.
+Model Strategy GSM8K MATH HumanEval MBPP MT-Bench
+LLaMA 2-7B Full FT 49.13±0.21 7.29±0.22 21.20±0.30 35.59±0.25 4.91±0.01
+LoRA(gaussian) 42.85±0.12 5.50±0.33 18.35±0.31 35.50±0.14 4.59±0.07
+LoRA(kaiming) 43.23±0.64 5.90±0.16 18.21±0.45 35.47±0.37 4.56±0.04
+PiSSA 53.22±0.55 7.47±0.34 21.92±0.38 37.24±0.63 4.88±0.03
+Mistral-7B Full FT 69.91±0.25 18.64±0.35 45.31±0.14 51.46±0.13 4.95±0.05
+LoRA(gaussian) 69.50±0.42 20.08±0.20 43.78±0.34 58.46±0.37 4.90±0.05
+LoRA(kaiming) 69.40±0.25 19.99±0.44 43.74±0.14 58.39±0.02 4.93±0.05
+PiSSA 73.31±0.23 23.12±0.52 46.88±0.25 62.55±0.58 5.34±0.04
+Gemma-7B Full FT 72.09±0.32 22.71±0.34 47.02±0.27 55.67±0.50 5.40±0.12
+LoRA(gaussian) 75.11±0.64 30.41±0.48 53.70±0.23 65.58±0.29 4.98±0.02
+LoRA(kaiming) 74.53±0.47 29.90±0.16 53.57±0.27 65.25±0.29 4.97±0.09
+PiSSA 77.78±0.32 31.33±0.33 54.31±0.28 66.17±0.43 5.64±0.10
 As shown in Table 2, across all models and tasks, fine-tuning with PiSSA consistently surpasses the performance of fine-tuning with LoRA.
-For instance, fine-tuning the LLaMA, Mistral, and Gemma with PiSSA results in performance improvements of 10.77%, 5.26%, and 3.04% respectively in mathematical tasks.
-In coding tasks, the improvements are 20%, 6.95%, and 1.14% respectively.
-In MT-Bench, we observe enhancements of 6.33%, 8.98%, and 13.25%.
-Notably, using PiSSA with only 2.3% of Gemma’s trainable parameters outperforms full parameter fine-tuning by 15.59% in coding tasks.
 Further experiments demonstrated that this improvement is robust across various amounts of training data and epochs (Section 5.2), including both 4-bit and full precision (Section 5.3), different model sizes and types (Section 5.4), and varying proportions of trainable parameters (Section 5.5).
-Table 2: Comparing PiSSA and LoRA on NLU tasks. Details can be found in Appendix  I.
-Method Parameters MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B
-RoBERTa-large (355M)
-Full FT 355M 90.2 96.4 90.9 68.0 94.7 92.2 86.6 91.5
-LoRA 1.84M 90.6 96.2 90.9 68.2 94.9 91.6 87.4 92.6
-PiSSA 1.84M 90.7 96.7 91.9 69.0 95.1 91.6 91.0 92.9
-DeBERTa-v3-base (184M)
-Full FT 184M 89.90 95.63 89.46 69.19 94.03 92.40 83.75 91.60
-LoRA 1.33M 90.65 94.95 89.95 69.82 93.87 91.99 85.20 91.60
-PiSSA 1.33M 90.43 95.87 91.67 72.64 94.29 92.26 87.00 91.88
-We also evaluate PiSSA’s natural language understanding (NLU) capability on the GLUE benchmark [61] with RoBERTa-large [62] and DeBERTa-v3-base [63].
-Table 2 presents the results across 8 tasks conducted with the two base models.
-PiSSA outperforms LoRA on 14 out of 16 experimental settings, and achieves the same performance on QQP + RoBERTa.
-The single exception arises from the DeBERTa-based model on the MNLI dataset.
-Upon reviewing the training loss, we observed that PiSSA’s average loss of $0.17$ was lower than LoRA’s $0.24$ in the final epoch.
+Table 3: Comparison of PiSSA and LoRA on NLU tasks.
+LoRAG and LoRAK denote LoRA with Gaussian and Kaiming initialization for $B$, respectively.
+Results for full fine-tuning, BitFit zaken2021bitfit , HAdapter houlsby2019parameter , PAdapter pfeiffer2020adapterfusion , LoRAG hu2021lora and AdaLoRA are from AdaLoRA zhang2023adaptive , averaged over five runs.
+Remaining methods are averaged over three runs, with details in Appendix L.
+Method Params MNLI SST2 MRPC CoLA QNLI QQP RTE STSB ALL
+Full FT 184M 89.90 95.63 89.46 69.19 94.03 92.40 83.75 91.60 88.25
+BitFit 0.1M 89.37 94.84 87.75 66.96 92.24 88.41 78.70 91.35 86.20
+HAdapter 1.22M 90.13 95.53 89.95 68.64 94.11 91.91 84.48 91.48 88.28
+PAdapter 1.18M 90.33 95.61 89.46 68.77 94.29 92.04 85.20 91.54 88.41
+LoRAG 1.33M 90.65 94.95 89.95 69.82 93.87 91.99 85.20 91.60 88.50
+LoRAK 1.33M 89.96 95.64 90.28 70.69 93.84 92.03 84.84 91.68 88.62
+DoRA 1.27M 90.29 95.79 90.93 70.85 94.10 92.07 86.04 91.79 88.98
+AdaLoRA 1.27M 90.76 96.10 90.69 71.45 94.55 92.23 88.09 91.84 89.46
+PiSSA 1.33M 90.37 96.22 91.50 73.12 94.43 92.33 88.69 92.00 89.83
+We also evaluate PiSSA’s natural language understanding (NLU) capability on the GLUE benchmark wang2018glue with DeBERTa-v3-base he2021debertav3 .
+Table 3 presents the results across 8 tasks.
+PiSSA outperforms LoRA in 7 out of 8 NLU tasks, achieving an overall average improvement of 1.21%.
+Upon reviewing the training loss on the exceptional dataset, MNLI, we observed that PiSSA’s average loss of $0.17$ was lower than LoRA’s $0.24$ in the final epoch.
 This indicates that the fitting ability of PiSSA remains stronger than that of LoRA.
@@
 The results are depicted in Figure 4.
-Additionally, similar comparisons on Mistral-7B and Gemma-7B are detailed in Appendix G.
+Additionally, similar comparisons on Mistral-7B and Gemma-7B are detailed in Appendix J.
 (a) Loss over steps.
@@
 (c) Accuracy on GSM8K over training steps.
-Figure 4: The loss, grad norm, and evaluation accuracy over the training steps of LoRA (indicated in blue), PiSSA (in orange), and full parameter fine-tuning (in green).
+Figure 4: The loss, grad norm, and evaluation accuracy over the training steps of LoRA (indicated in blue), PiSSA (in orange), and full parameter fine-tuning (in red).
 According to Figure 4(a), the loss of PiSSA reduces rapidly during the first 100 steps, and the grad norm (shown in Figure 4(b)) of PiSSA is significantly higher than that of LoRA, with a trend similar to full fine-tuning.
@@
 The reduction is more significant for lower-rank matrices.
-For instance, in the LLaMA-3-70B [64], all “Key” projection layers see a reduction of 49%.
+For instance, in the LLaMA-3-70B dubey2024llama , all “Key” projection layers see a reduction of 49%.
 The results in Table 4 validate that QLoRA, discussed in Section 4, does not reduce quantization error.
-In contrast, PiSSA significantly outperforms LoftQ in reducing quantization error, as further discussed in Appendix F.
-Besides reducing the quantization error, we expect QPiSSA to also converge faster than QLoRA and LoftQ.
-We train LLaMA 3-8B with LoRA/QLoRA, PiSSA/QPiSSA, LoftQ, and full fine-tuning on MetaMathQA-395K for 3 epochs and record the loss, grad norm, and accuracy on GSM8K.
+In contrast, PiSSA significantly outperforms LoftQ in reducing quantization error, as further discussed in Appendix H.
 (a) Loss over training steps.
@@
 Figure 5: The loss, grad norm, and evaluation accuracy over the training steps of (Q)LoRA, (Q)PiSSA, LoftQ and full parameter fine-tuning.
+The difference between QPiSSA and PiSSA is the quantization of the residual model to 4 bits.
+As introduced in Section 4, the residual model has less influence on the optimal direction compared with the principal adapter, which is the same in both QPiSSA and PiSSA.
+Therefore, besides reducing the quantization error, we expect QPiSSA to also converge faster than QLoRA and LoftQ.
+We train LLaMA 3-8B using LoRA/QLoRA, PiSSA/QPiSSA, LoftQ, and full fine-tuning on MetaMathQA-395K for 3 epochs, recording the loss, gradient norm, and accuracy on GSM8K, as shown in Figure 5.
 According to Figure 5, QPiSSA’s loss reduction speed in the first 100 steps is even faster than PiSSA and full fine-tuning.
@@
 5.4 Experiments Across Various Sizes and Types of Models
-In this section, we compare (Q)PiSSA and (Q)LoRA across 9 models, ranging from 7-70B parameters, including LLaMA 2-7/13B [54], LLaMA-3-8/70B [64], Mistral-7B [55], Gemma-7B [56], and Qwen1.5-7B [65], Yi-1.5-34B [66] and MoE models: DeepSeek-MoE-16B [67] and Mixtral-8x7B [68].
+In this section, we compare (Q)PiSSA and (Q)LoRA across 9 models, ranging from 7-70B parameters, including LLaMA 2-7/13B touvron2023llama , LLaMA-3-8/70B dubey2024llama , Mistral-7B jiang2023mistral , Gemma-7B team2024gemma , and Qwen1.5-7B qwen , Yi-1.5-34B young2024yi and MoE models: DeepSeek-MoE-16B dai2024deepseekmoe and Mixtral-8x7B jiang2024mixtral .
 These models were fine-tuned on the MetaMathQA-100K and CodeFeedback-100K dataset and evaluated on the GSM8K and HumanEval.
 DeepSeek-MoE-16B, Mixtral-8x7B, Yi-1.5-34B, and LLaMA-3-70B were fine-tuned with QPiSSA and QLoRA, while the other models were using PiSSA and LoRA.
-From Figure 6, it can be observed that (Q)PiSSA, compared to (Q)LoRA, shows improved accuracy across various sizes and types of models, demonstrating its consistent advantage over (Q)LoRA.
+From Figure 6, (Q)PiSSA, compared to (Q)LoRA, shows improved accuracy across various sizes and types of models, demonstrating its consistent advantage over (Q)LoRA.
 Figure 6: Comparison of (Q)PiSSA and (Q)LoRA across models from 7B to 70B.
@@
 The training is conducted using the MetaMathQA-100K dataset for 1 epoch, while the validation is performed on the GSM8K and MATH datasets.
-The outcomes of these experiments are depicted in Figure 7, with additional results presented in Appendix H.
+The outcomes of these experiments are depicted in Figure 7, with additional results presented in Appendix K.
 (a) Quantization error reduction ratio across ranks.
@@
 This paper presents a PEFT technique that applies singular value decomposition (SVD) to the weight matrix of pre-trained models.
-The principal components obtained from the SVD are used to initialize a low-rank adapter named PiSSA, while the residual components are kept frozen, in order to achieve effective fine-tuning and parameter efficiency at the same time.
+The principal components obtained from the SVD are used to initialize a low-rank adapter named PiSSA, while the residual components are kept frozen, to achieve effective fine-tuning and parameter efficiency simultaneously.
 Through extensive experiments, we found that PiSSA and its 4-bit quantization version QPiSSA significantly outperform LoRA and QLoRA in both NLG and NLU tasks, across different training steps, various model sizes and types, and under various amount of trainable parameters.
@@
 There are still some questions with PiSSA not addressed in this paper: 1) Besides language models, can PiSSA also be adapted to convolutional layers and enhance the performance of vision tasks?
-2) Can PiSSA also benefit from some improvements to LoRA, such as AdaLoRA [69] and DyLoRA [42] which adaptively adjust the rank?
+2) Can PiSSA also benefit from some improvements to LoRA, such as AdaLoRA zhang2023adaptive and DyLoRA valipour2022dylora which adaptively adjust the rank?
 3) Can we provide more theoretical explanations for the advantages of PiSSA over LoRA?
@@
 Nevertheless, we are excited to see the huge potential of PiSSA already demonstrated in existing experiments and look forward to more tests and suggestions from the community.
-Appendix A Conductive Experiments on Various SVD Components
-To investigate the influence of singular values and vectors of varying magnitudes on the fine-tuning performance, we initialize the adapters injected into LLaMA 2-7B, Mistral-7B-v0.1, and Gemma-7B with principal, medium, and minor singular values and vectors.
-These models are then fine-tuned on the MetaMathQA dataset [2] and evaluated against the GSM8K [57] and MATH datasets [70], with the outcomes depicted in Figures 8.
-Figure 8: Initializing with principal, medium, and minor singular values and vectors, the training loss on the MetaMathQA and the accuracy on the GSM8K and MATH validation sets are reported, respectively, for three models.
-The results highlight that initializing adapters with principal singular values and vectors consistently leads to reduced training loss and enhanced accuracy on both the GSM8K and MATH validation datasets across all three models.
-This underscores the efficacy of our strategy in fine-tuning the model parameters based on the principal singular values.
+8 Acknowledgements:
+This work is supported by the National Key R&D Program of China (2022ZD0160300).
+Appendix A Enhancing PiSSA with LoRA Improvement Methods
+AdaLoRA introduces three improvements over LoRA:
+  • Trainable parameters in AdaLoRA are changed to $A,B$, and $E$.
+$A$ and $B$ are Gaussian-initialized, and $E$ is a zero-initialized $r$-dimensional vector, making $Adiag(E)B=\Delta W$, similar to singular value decomposition.
+  • A regularization loss $|AA^{T}-I|+|B^{T}B-I|$ is used to make $A$ and $B$ gradually orthogonal during training, resembling the SVD of $\Delta W$.
+  • An initial large rank is set, and less important E values are gradually masked during training, resulting in different final ranks for each layer, achieving better performance with the same number of parameters.
+Despite the extensive use of SVD terms, AdaLoRA does not perform actual SVD on any matrix.
+In the PEFT domain, terms like low-rank decomposition, and singular value decomposition often appear.
+They generally refer to products of low-dimensional matrices approximating an ideal $\Delta W$ without actual matrix decomposition.
+To our knowledge, PiSSA is the first to perform SVD on the original model, fine-tuning the principal component while keeping the residual model frozen.
+PiSSA and AdaLoRA represent different improvements to LoRA, making them combinable. Therefore, we additionally improved PiSSA based on AdaLoRA’s three innovations:
+  • After extracting the principal singular values and vectors of $W$, we use $S$ as an independent trainable vector instead of multiplying it into $U$ and $V$.
+  • Since PiSSA’s $U$ and $V$ are orthogonal at the beginning, maintaining their orthogonality through orthogonal regularization is very easy.
+  • Although AdaLoRA claims to dynamically reduce the number of trainable parameters, the initially large number of parameters is not truly pruned, resulting in more parameters being updated during actual training.
+Therefore, we did not use this improvement.
+DoRA adds a learnable magnitude module to LoRA, normalizing $W+AB$ at each update step and multiplying its by the magnitude module.
+This allows $A,B$ to learn the direction and the magnitude module to learn the magnitude of $\Delta W$.
+While this approach can improve fine-tuning performance, normalizing $W+AB$ at each step results in slower fine-tuning speeds.
+In contrast, PiSSA only changes LoRA’s initialization method, matching LoRA in training speed and converging faster, thereby reducing training costs.
+Table 5: GSM8K accuracy for LoRA and PiSSA when combined with LoRA improvement methods.
+Model Method LoRA+ PiSSA+
+Vanilla 71.01±0.199 76.75±0.036
+LLaMA-3-8B DoRA 72.38±0.189 77.51±0.257
+AdaLoRA 72.31±0.202 78.59±0.199
+PiSSA, with its intrinsic principal singular values and orthogonal singular vectors, is very suitable for combination with AdaLoRA.
+According to Table 5.
+The performance of the improved PiSSA surpasses all the other methods including PiSSA.
+From lines 1, and 2 of the table, it is evident that the performance of PiSSA combined with DoRA significantly surpasses that of DoRA alone and also exceeds the performance of PiSSA alone.
+Taking into account the combination experiments of PiSSA with AdaLoRA, it can be inferred that PiSSA benefits from the enhancement techniques of LoRA, demonstrating the potential of PiSSA when integrated with other methods.
 Appendix B Fast Singular Value Decomposition
@@
 For QLoRA, QPiSSA and LoftQ, the base model was used NF4 precision, while the adapter layers used FP32 precision.
-Appendix E Reducing Quantization Error through QPiSSA with Multiple Iteration of SVD
+Appendix E Reducing Quantization Error through Multiple Iteration of SVD
 Table 8 provides a supplementary explanation of the results in Table 4.
@@
 In the main paper, we used 5 iterations in Section 5.3 and Section 5.4, while 1 iteration was used in Section 5.5.
-Appendix F Comparing the Quantization Error of QLoRA, LoftQ and QPiSSA
+Appendix F Conductive Experiments on Various SVD Components
+To investigate the influence of singular values and vectors of varying magnitudes on the fine-tuning performance, we initialize the adapters injected into LLaMA 2-7B, Mistral-7B-v0.1, and Gemma-7B with principal, medium, and minor singular values and vectors.
+These models are then fine-tuned on the MetaMathQA dataset [2] and evaluated against the GSM8K [54] and MATH datasets [55], with the outcomes depicted in Figures 8.
+Figure 8: Initializing with principal, medium, and minor singular values and vectors, the training loss on the MetaMathQA and the accuracy on the GSM8K and MATH validation sets are reported, respectively, for three models.
+The results highlight that initializing adapters with principal singular values and vectors consistently leads to reduced training loss and enhanced accuracy on both the GSM8K and MATH validation datasets across all three models.
+This underscores the efficacy of our strategy in fine-tuning the model parameters based on the principal singular values.
+Appendix G The Residual Matrices having a Narrower Distribution
+To intuitively compare the distribution differences between quantized original and residual models, in Figure 3, we took LLaMA 2-7B’s first Query layer as an example to illustrate the distribution of $W$ and $W_{res}$.
+However, using only one layer of one model is not statistically significant.
+In this section, we applied PiSSA initialization to LLaMA-2-7B, Mistral-7B, Gemma-7B, and LLaMA-3-8B, and fit the values in every linear layer with Gaussian distribution and calculated their mu and sigma.
+(a) LLaMA-2-7B
+(b) Mistral-7B-v0.1
+(c) Gemma-7B
+(d) LLaMA-3-8B
+Figure 9: Comparison of Loss and Ratio to the target A and target B for LoRA and PiSSA across the initial 5 steps.
+The results in Figure 9 show that the residual models’ means are closer to 0, and the standard deviations are smaller after PiSSA initialization.
+Thus, $W^{res}$ indeed has a narrower distribution than W in a statistical sense.
+Nevertheless, the difference is not as large as that in the first layer after averaging all layers, which we suspect is because middle layers in a model tend to have more even eigenvalue distributions due to redundancy and insufficient training.
+Appendix H Comparing the Quantization Error of QLoRA, LoftQ and QPiSSA
 This section extends the discussion in Section 4 by providing a comprehensive comparison of the quantization errors associated with QLoRA, LoftQ, and QPiSSA.
@@
 Figure 11: Fitting the original matrix and the residual matrix using Student’s t-distribution.
-Appendix G Evaluating PiSSA on Mixtral and Gemma with More Training Steps
+Appendix I Combining QPiSSA with Various Quantization Methods
+In addition to NF4 quantization, QPiSSA can also be combined with GPTQ and INT8 quantization. We posit that PiSSA effectively reduces quantization error for several reasons:
+  • It reduces outlier values;
+  • It makes the value distribution more Gaussian-like;
+  • It preserves larger values in full precision, thereby narrowing the weight distribution in the quantized portion.
+While INT8 also targets the reduction of outlier values (point 1), PiSSA has the potential to enhance this effect.
+The second point aligns well with NF4, and the third point is crucial as PiSSA uses an adaptor to retain a significant portion of weights in full precision, maintaining the integrity of critical values.
+Table 9: Quantization Error and Accuracy for PiSSA Combined with Various Quantization Methods.
+GPTQ quantizes each row $w$ independently, adjusting one weight at a time while updating all remaining, non-quantized weights.
+Therefore, the nuclear norm method used for calculating quantization error in the main paper is not applicable to GPTQ.
+Instead, we measure Perplexity on WikiText-2, where a lower Perplexity indicates reduced quantization error.
+Model Dataset Quantization Error GSM8K Accuracy
+QLoRA PiSSA QLoRA PiSSA
+NF4 324.8 (nuclear norm) 265.8 (nuclear norm) 70.79±0.42 73.76±0.20
+LLaMA-3-8B INT8 34.47 (nuclear norm) 28.21 (nuclear norm) 71.68±0.14 76.54±0.32
+GPTQ 20.79 (perplexity) 6.23 (perplexity) 70.18±0.42 74.58±0.22
+As shown in Table 9, QPiSSA combined with INT8 reduces quantization error by 18.16% on LLaMA-3-8B, and significantly outperforms QLoRA using INT8.
+Furthermore, in row 3 of the table, the perplexity of LLaMA-3-8B increases to 20.79 after quantization with GPTQ-4bit on the C4 dataset.
+However, when PiSSA is applied, the perplexity is reduced to 6.23.
+These results confirm the effectiveness of PiSSA in reducing quantization error, as discussed in the main paper.
+Overall, QPiSSA demonstrates a clear advantage over QLoRA when combined with various quantization methods, retaining the fast convergence and superior performance characteristics of PiSSA while minimizing quantization error.
+Appendix J Evaluating PiSSA on Mixtral and Gemma with More Training Steps
 This is the supplement for Section 5.2.
 We applied PiSSA, LoRA, and full parameter fine-tuning on the full MetaMathQA-395K dataset using Mistral-7B and Gemma-7B models, training for 3 epochs.
-Figures 11 and 12 display the training loss, gradient norm, and evaluation accuracy on GSM8K.
+Figures 12 and 13 display the training loss, gradient norm, and evaluation accuracy on GSM8K.
 (a) Loss over training steps.
@@
 Figure 13: Fine-tuning Gemma-7B on the MetaMathQA-395K dataset for 3 epochs: A comparison of full parameter fine-tuning (indicated by a dashed line), LoRA (in blue), and PiSSA (in orange).
-As shown in Figure 10(a) and 11(a), the loss for full parameter fine-tuning decreases sharply with each epoch, indicating overfitting to the training data.
+As shown in Figure 11(a) and 12(a), the loss for full parameter fine-tuning decreases sharply with each epoch, indicating overfitting to the training data.
 Notably, during the entire first epoch, the loss for full parameter fine-tuning on Mistral and Gemma is significantly higher than for LoRA and PiSSA, suggesting that full parameter fine-tuning has weaker generalization capabilities compared to LoRA and PiSSA on Mistral-7B and Gemma-7B models.
 The gradient norm for the first epoch in Figure 12(b) fluctuates dramatically with each step, further indicating instability in the training process for full parameter fine-tuning.
-Consequently, as illustrated in Figures 10(c) and 11(c), the performance of full parameter fine-tuning is markedly inferior to that of LoRA and PiSSA.
+Consequently, as illustrated in Figures 11(c) and 12(c), the performance of full parameter fine-tuning is markedly inferior to that of LoRA and PiSSA.
 These experiments demonstrate that using parameter-efficient fine-tuning can prevent the over-fitting issue caused by over-parameters.
-Appendix H Supplementary Experiments on Various Ranks
-H.1 Quantization Error for More Type of Layers
+Appendix K Supplementary Experiments on Various Ranks
+K.1 Quantization Error for More Type of Layers
 Figure 7(a) only shows the reduction ratio of quantization error for “q_proj” layers.
@@
 From Figure 14 it can be observed that under different ranks, the reduction ratio of quantization error for various linear layers in LLaMA-2-7B, including “k_proj”, “v_proj”, “o_proj”, “gate_proj”, “up_proj”, and “down_proj” layers, is consistently lower with PiSSA compared to LotfQ.
-H.2 Evaluation Performance for More Model on Various Ranks
+K.2 Evaluation Performance for More Model on Various Ranks
 Section 5.5 only validated the effectiveness of LLaMA-2-7B. In Figure 15, we also present the comparative results of Mistral-7B-v0.1, and Gemma-7B under different ranks.
@@
 This over-parameterization phenomenon does not occur on LLaMA-2-7B, suggesting that increasing the rank further might enable PiSSA to achieve even higher performance on LLaMA-2-7B.
-H.3 More Training Loss and Grad Norm under Various Ranks
-In Figure 15 and 16, we examining the loss and gradient norm during the training process of PiSSA and LoRA on LLaMA 2-7B, Mistral-7B-v0.1, and Gemma-7B using different ranks.
+K.3 More Training Loss and Grad Norm under Various Ranks
+In Figure 16 and 17, we examining the loss and gradient norm during the training process of PiSSA and LoRA on LLaMA 2-7B, Mistral-7B-v0.1, and Gemma-7B using different ranks.
 Figure 16: Comparison of training loss for LLaMA-2-7B, Mistral-7B, and Gemma-7B, organized into three rows, using LoRA and PISSA across ranks $2^{i},i\in[0,7]$, organized into eight columns.
@@
 It demonstrates the robustness of the faster convergence property of PiSSA across various ranks.
-Appendix I Experimental Settings on NLU
+Appendix L Experimental Settings on NLU
 Datasets
@@
 Implementation Details
-To evaluate the performance of PiSSA intuitively, we compared PiSSA and LoRA with the same number of trainable parameters.
-RoBERTa-large has $355M$ trainable parameters.
-PiSSA was applied to $W_{Q}$ and $W_{V}$, resulting in a total of $1.84M$ trainable parameters.
+To evaluate the performance of PiSSA intuitively, we compared PiSSA to LoRA with the same number of trainable parameters.
 DeBERTa-v3-base has $184M$ trainable parameters.
 PiSSA and LoRA were applied to $W_{Q}$, $W_{K}$ and $W_{V}$ respectively, resulting in a total of $1.33M$ trainable parameters.
-Experiments on NLU are based on the publicly available LoftQ [49] code-base.
-On RoBERTa-large, we initialize the model with pretrained MNLI checkpoint for the task of MRPC, RTE and STS-B.
+The results for full fine-tune, BitFit [15], HAdapter [30], PAdapter [36], LoRA with Gassian initialization [11] and AdaLoRA are sourced from AdaLoRA [59], based on five runs.
+The remaining results use the publicly available LoftQ [14] codebase and are averaged over three runs.
+In LoRA, the B matrix is initialized to zero, while the A matrix can be initialized using various methods, such as Gaussian initialization and Kaiming initialization [68].
+The selection of the initialization method can influence the final results.
+In this paper, we report the different results of LoRA based on Gaussian initialization and Kaiming initialization in the experiments, as shown in Table 2 and Table 3.
+For DoRA, we used the code from the PEFT package for deployment and conducted a search on key hyperparameters.
 We set the rank of PiSSA in this experiment as $8$ and selecte lora alpha in 8, 16.
@@
 Batch sizes (BS) are selected from ${6,8,16,32}$.
-Table 7 presents the hyperparmeters we utilized on the glue benchmark.
-Table 7: Hyperparameters of our Methods on GLUE.
-Dataset Roberta-large Deberta-v3-base
-Epoch BS LR Lora alpha Epoch BS LR Lora alpha
-MNLI 10 32 1e-4 16 5 16 5e-5 8
-SST-2 10 32 2e-4 16 20 16 3e-5 8
-MRPC 20 16 6e-4 8 20 32 2e-4 8
-CoLA 20 16 4e-4 8 20 16 1e-4 8
-QNLI 10 6 1e-4 8 10 32 1e-4 16
-QQP 20 32 3e-4 8 10 16 1e-4 8
-RTE 20 16 3e-4 16 50 16 1e-4 8
-STS-B 30 16 3e-4 16 20 8 3e-4 8
+The hyperparameter configurations of PiSSA, DoRA and LoRA with Kaiming Initialization are shown in Table 10.
+LoRAK denotes LoRA with Kaiming initialization, and $\alpha$ denotes LoRA alpha.
+Table 10: Hyperparameters of PiSSA, DoRA and LoRA with Kaiming Initialization on GLUE.
+Dataset PiSSA DoRA LoRAK
+Epoch BS LR $\alpha$ Epoch BS LR $\alpha$ Epoch BS LR $\alpha$
+MNLI 5 16 5e-4 8 10 32 2e-4 16 10 32 3e-4 8
+SST-2 20 16 3e-5 8 10 16 4e-4 16 10 32 1e-4 8
+MRPC 20 32 2e-4 8 10 32 4e-4 16 10 32 4e-4 8
+CoLA 20 16 1e-4 8 20 8 1e-4 6 30 32 4e-4 8
+QNLI 10 32 1e-4 16 10 16 2e-4 16 25 32 3e-4 8
+QQP 10 16 1e-4 8 10 16 1e-4 6 10 16 3e-4 8
+RTE 50 16 1e-4 8 50 8 2e-4 6 50 32 4e-4 8
+STS-B 20 8 3e-4 8 20 16 3e-4 6 30 16 4e-4 8
+Appendix M Comparison of Initial Gradient Subspaces
+To compare the gradient subspaces of PiSSA and LoRA, we conducted two additional experiments to validate our analysis.
+First, we trained LLaMA-3-8B on the MetaMath dataset five times, initializing LoRA with different random seeds while using the same batch of 128 training examples to compute LoRA’s gradients.
+After performing dimensionality reduction to two dimensions, the results are presented in Table 11.
+Table 11: Ablation results for LoRA and PiSSA across different seeds.
+Method Seed 0 Seed 1 Seed 2 Seed 3 Seed 4
+grad_A LoRA [0,0] [0,0] [0,0] [0,0] [0,0]
+PiSSA [0,1] [0,1] [0,1] [0,1] [0,1]
+grad_B LoRA [-0.99, 0.12] [0.95, 0.31] [0.46, -0.89] [0.24, 0.97] [0.04, -0.99]
+PiSSA [1,0] [1,0] [1,0] [1,0] [1,0]
+We observe that the gradient of matrix $A$ remains consistently zero, while the gradient direction of matrix $B$ varies across initializations.
+This behavior arises because matrix $A$‘s gradient depends on matrix $B$, which in LoRA is initialized to zero, resulting in a zero gradient for $A$.
+In contrast, matrix $B$ is initialized from a Gaussian distribution, leading to variation in its gradient direction across different seeds.
+In comparison, PiSSA’s gradient direction remains consistent across all five training runs, as it solely depends on the original model and the training data.
+This experiment highlights the stability of PiSSA’s optimization trajectory relative to LoRA’s more variable directionality.
+Next, we quantitatively compared the effect of updating along the principal singular value direction versus a “random” direction during the early stages of fine-tuning.
+We trained LLaMA-3-8B on the MetaMathQA dataset using both PiSSA and LoRA, saving the parameters and gradients from the first 50 iterations.
+At the 50th step, the loss values for LoRA and PiSSA were 0.3677 and 0.2899, respectively.
+Using the parameters from the 50th step as the target point, we evaluated the movement in the first five steps relative to the target, computing how much progress was made towards the final point.
+We then divided this progress by the total target distance to obtain a ratio.
+These ratios are shown in Figure 18.
+(a) Loss over steps.
+(b) Ratio to target A over steps.
+(c) Ratio to target A over steps.
+Figure 18: Comparison of Loss and Ratio to the target A and target B for LoRA and PiSSA across the initial 5 steps.
+The results reveal that after just five updates, PiSSA reduced the loss from 0.8884 to 0.3346, while LoRA’s loss reduction was more modest, dropping to only 0.5538.
+This demonstrates the advantage of updating along the principal singular value direction, which PiSSA leverages, leading to faster convergence.
+Further, in the first step, matrix $A$ in LoRA exhibited a zero gradient and therefore did not update.
+Over the next four steps, it moved only 15.94% towards the target direction.
+Similarly, matrix $B$ in LoRA consistently moved less towards the target endpoint compared to PiSSA.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: 1
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: 7
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory Assumptions and Proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: 3, 4
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental Result Reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: 5
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: 5
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental Setting/Details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: 5
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment Statistical Significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [No]
+  34. Justification: error bars are not reported because it would be too computationally expensive
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments Compute Resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: 5
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code Of Ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: the paper conform with the NeurIPS Code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader Impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [N/A]
+  49. Justification: We only use public available datasets.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: We only use public available datasets and models.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: 2
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New Assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification: the paper does not release new assets.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and Research with Human Subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: the paper does not involve crowdsourcing nor research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: the paper does not involve crowdsourcing nor research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.